### v1

# Frobenius intertwiners for $q$-difference equations

Abstract.
In this note we consider a class of $q$-hypergeometric equations describing the quantum difference equation for $X=T^{*}\mathbb{P}^{1}$ We show that over $\mathbb{Q}_{p}$ these equations are equipped with the Frobenius action $z\to z^{p}$. We obtain an explicit formula for the constant term of the Frobenius intertwiner in therms of $p$ adic $q$ gamma function of Koblitz. In the limit $q\to 1$ our formula degenerates to the Frobenius intertwiners for the hypergeometric differential equations discovered by B.Dwork.
Department of Mathematics, University of North Carolina at Chapel Hill
Chapel Hill, NC 27599-3250, USA
Key words: Frobenius intertwiners, $q$ hypergeometric functions, $p$ adic $q$ Gamma function
2020 Mathematics Subject Classification: 14G33 (11D79, 32G34, 33C05, 33E30)
footnotetext: ${}^{\star}\mskip-0.99998mu$-mail: asmirnov@email.unc.edu
1. q-hypergeometric functions
In this section we let $\mathbb{K}=\mathbb{C}$. In the next section we switch to the field of $p$-adic numbers.
1.1.
Let $D$ be the $q$-shift operator acting by $Df(z)=f(zq)$. We consider a $q$-holonomic module $\mathcal{H}({\boldsymbol{u}},\hbar,q,z)$ with $\mathbb{K}^{\times}$-valued parameters ${\boldsymbol{u}}=(u_{1},\dots,u_{n})$, $\hbar$ given by the following $q$-difference equation
(1.1) $\Big{(}\prod\limits_{i=1}^{n}\dfrac{1-u^{-1}_{i}D}{1-\hbar q^{-1}u^{-1}_{i}D}\Big{)}F(z)=zF(z).$
This $q$-difference equation can be written in a more standard form:
(1.2) $P(\boldsymbol{a},\hbar,z)F(z)=0,$
where $P(\boldsymbol{a},\hbar,z)$ denotes the $q$-difference operator:
(1.3) $P(\boldsymbol{a},\hbar,z)=\sum_{m=0}^{n}\,\alpha_{m}(1-z\hbar^{m})D^{m}.$
and $\alpha_{m}$ denotes the coefficient of $x^{m}$ in $\prod\limits_{i=1}^{n}(1-x/u_{i})$.
1.2.
Let us consider the $q$-hypergeometric power series
(1.4) $F_{i}({\boldsymbol{u}},\hbar,q,z)=\sum\limits_{d=0}^{\infty}\,\Big{(}\prod\limits_{j=1}^{n}\,\dfrac{(u_{i}\hbar/u_{j})_{d}}{(u_{i}q/u_{j})_{d}}\Big{)}\,z^{d}.$
Set $e_{i}(z)=\exp\Big{(}{\frac{\ln(z)\ln(u_{i})}{\ln(q)}}\Big{)},$ so that $e_{i}(zq)=e_{i}(z)u_{i}$. One verifies that the functions
(1.5) $\tilde{F}_{i}({\boldsymbol{u}},\hbar,q,z)=e_{i}(z)F_{i}({\boldsymbol{u}},\hbar,q,z)$
satisfy (1.1) for all $i=1,\dots,n$. Thus, $\tilde{F}_{i}({\boldsymbol{u}},\hbar,q,z)$ give a basis of solutions for $\mathcal{H}({\boldsymbol{u}},\hbar,q,z)$.
1.3.
From (1.2) it is clear that the vectors
(1.10) ${\Psi}_{i}({\boldsymbol{u}},\hbar,q,z)=\left(\begin{array}[]{l}\tilde{F}_{i}({\boldsymbol{u}},\hbar,q,z)\\ \tilde{F}_{i}({\boldsymbol{u}},\hbar,q,zq)\\ \dots\\ \tilde{F}_{i}({\boldsymbol{u}},\hbar,q,zq^{n-1})\end{array}\right)$
satisfy a first order $q$-difference equation
(1.11) ${\Psi}_{i}({\boldsymbol{u}},\hbar,q,zq)=\mathbf{M}_{\mathcal{L}}({\boldsymbol{u}},\hbar,q,z)\,{\Psi}_{i}({\boldsymbol{u}},\hbar,q,z)$
for all $i=1,\dots,n$, where $\mathbf{M}_{\mathcal{L}}({\boldsymbol{u}},\hbar,q,z)$ is the companion matrix for the characteristic polynomial of (1.3):
$\mathbf{M}_{\mathcal{L}}({\boldsymbol{u}},\hbar,q,z)=\left(\begin{array}[]{cccccc}0&1&0&0&\dots&0\\ 0&0&1&0&\dots&0\\ 0&0&0&1&\dots&0\\ ..&..&..&..&..&..\\ \frac{\bar{\alpha}_{n}(z-1)}{1-\hbar^{n}z}&\frac{\bar{\alpha}_{n-1}(z\hbar-1)}{1-\hbar^{n}z}&\frac{\bar{\alpha}_{n-2}(z\hbar^{2}-1)}{\hbar^{n}z-1}&\frac{\bar{\alpha}_{n-3}(z\hbar^{3}-1)}{1-\hbar^{n}z}&\dots&\frac{\bar{\alpha}_{1}(z\hbar^{n-1}-1)}{1-\hbar^{n}z}\\ \end{array}\right)$
where $\bar{\alpha}_{k}$ denotes the coefficients of $x^{k}$ in $\prod_{i=1}^{n}(1-xu_{i})$. The matrix
${\Psi}({\boldsymbol{u}},\hbar,q,z)=\Big{(}{\Psi}_{1}({\boldsymbol{u}},\hbar,q,z),\dots,{\Psi}_{n}({\boldsymbol{u}},\hbar,q,z)\Big{)}.$
with $i$-th column given by ${\Psi}_{i}({\boldsymbol{u}},\hbar,q,z)$ is a fundamental solution matrix of the $q$-hypergeometric system $\mathcal{H}({\boldsymbol{u}},\hbar,q,z)$. Explicitly, it has the following matrix elements
(1.12) ${\Psi}_{i,j}({\boldsymbol{u}},\hbar,q,z)=\tilde{F}_{j}({\boldsymbol{u}},\hbar,q,zq^{i-1}).$
Taking a matrix with columns given by $\mathbb{K}$ linear combinations of the columns of $\tilde{\Psi}({\boldsymbol{u}},\hbar,q,z)$ produces new fundamental solution matrix. Thus other fundamental solution matrices are of the form
${\Psi}({\boldsymbol{u}},\hbar,q,z)\Lambda,\ \ \Lambda\in\mathsf{GL}_{n}(\mathsf{\mathbb{K}}).$
Finally, we also note that the fundamental solution matrix has the form
(1.13) ${\Psi}({\boldsymbol{u}},\hbar,q,z)={\Psi}^{\prime}({\boldsymbol{u}},\hbar,q,z)\textsf{E}(z).$
where $\textsf{E}(z)$ is the diagonal matrix $\textsf{E}(z)=\operatorname{diag}(e_{1}(z),e_{2}(z),\dots,e_{n}(z))$, and ${\Psi}^{\prime}({\boldsymbol{u}},\hbar,q,z)$ is analytic near $z=0$, i.e., is a power series in $z$.
1.4.
Throughout this paper we will be using both, the multiplicative parameters ${\boldsymbol{u}}=(u_{1},\dots,u_{n})$, $\hbar$ and additive parameters $\boldsymbol{a}=(a_{1},\dots,a_{n})$, $h$ which are related by
(1.14) $u_{1}=q^{a_{1}},\dots,u_{n}=q^{a_{n}},\ \ \hbar=q^{h}.$
2. Vertex of $X=T^{*}\mathbb{P}^{n-1}$
The $q$-hypergeometric system $\mathcal{H}({\boldsymbol{u}},\hbar,q,z)$ describes equivariant quantum $K$-theory of the cotangent bundles over projectile spaces. This set up will be the most convenient to us and we briefly describe it in this section.
2.1.
Let (×$(\mathbb{C}^{\times})^{n}$ be the torus acting on $\mathbb{C}^{n}$ by scaling the coordinate lines with characters $u_{1},\dots,u_{n}$. We have an induced action of this torus on the cotangent bundle over projectile space $X=T^{*}\mathbb{P}(\mathbb{C}^{n})$. Let $T=(\mathbb{C}^{\times})^{n}\times\mathbb{C}^{\times}$ be a bigger torus acting on $X$: the first factor (×$(\mathbb{C}^{\times})^{n}$ acts as before and the second $\mathbb{C}^{\times}$ scales the cotangent directions with a character $\hbar^{-1}$. Recall that the $T$-equivariant K-theory of $X$ is a ring:
(2.1) $K_{T}(X)=\mathbb{Z}[L^{\pm},u_{1}^{\pm},\dots,u^{\pm}_{n},\hbar^{\pm}]/(L-u_{1})\dots(L-u_{n}).$
where $L$ is the class of the line bundle $\mathcal{O}(1)$ and abusing notations we denote by the same symbols $u_{1},\dots,u_{n}$ the trivial equivariant line bundles associated to the characters $u_{i}$.
In this language, the additive parameters $a_{i}$, $h$ may be thought of as the first Chern classes of the line bundles $u_{i}$ and $\hbar$. In the same way, let $x$ be the first Chern class of $L$, i.e., we define $L=q^{x}$.
2.2.
Let us define a formal extension of this ring
$\hat{K}_{T}(X)=\mathrm{completion\ of}\ \ K_{T}(X)[z^{x}][[q]],$
so that it contains the following gamma class:
(2.2) $\Phi(\boldsymbol{a},h,x,q,z)=\Big{(}\dfrac{(q^{h},q)_{\infty}}{(q,q)_{\infty}}\Big{)}^{\!\!n}\,\prod\limits_{i=1}^{n}\dfrac{(q^{\,x+1-a_{i}},q)_{\infty}}{(q^{\,x+h-a_{i}},q)_{\infty}}\,z^{\,x}\in\hat{K}_{T}(X),$
where $(u,q)_{\infty}=\prod\limits_{m=0}^{\infty}\,(1-uq^{m})$ is the reciprocal of the $q$ gamma function. The normalized vertex function of $X$ is a generating functions counting equivariant quasimaps to quiver varieties, see Section 7.2 of [Oko17] for definitions. For $X$ this functions has the following form:
(2.3) $\widetilde{V}(\boldsymbol{a},x,h,q,z)=\sum\limits_{d=0}^{\infty}\,\Phi(\boldsymbol{a},h,x+d,q,z)\,\in\hat{K}_{T}(X)[[z]]$
One verifies that
$P({\boldsymbol{u}},\hbar,z)\,\widetilde{V}(\boldsymbol{a},x,h,q,z)=(L-u_{1})(L-u_{2})\dots(L-u_{n})\,\Phi(\boldsymbol{a},h,x,q,z)$
where $P({\boldsymbol{u}},\hbar,z)$ is the $q$-difference operator given by (1.3). This means that in (2.1) the vertex function solves the $q$-hypergeometric equation (1.2).
2.3.
The components of the class (2.3) in any chosen basis of $K$-theory (2.1) provide a basis of solutions of the $q$-hypergeometric system $\mathcal{H}({\boldsymbol{u}},\hbar,q,z)$. For example, one convenient basis consists of the $K$-theory classes of the $T$-fixed points $X^{T}$. The components of the class (2.3) in this basis are obtained by the substitution $L=u_{i}$ or, in the additive notations, by the substitution $x=a_{i}$. In this way we obtain the basis of solutions given, up to a normalization multiple, by the hypergeometric series (1.5):
(2.4) $\widetilde{V}(\boldsymbol{a},a_{i},h,q,z)=\Phi(\boldsymbol{a},h,a_{i},q,z)\tilde{F}_{i}(\boldsymbol{a},h,q,z).$
Therefore, the vertex function (2.3) provides a uniform “basis free” description of solutions for the $q$-hypergeometric system $\mathcal{H}({\boldsymbol{u}},\hbar,q,z)$.
2.4.
It will be convenient to normalize all solutions as in the vertex (2.4). In particular, we introduce the normalized fundamental solution matrix
$\tilde{\Psi}(\boldsymbol{a},h,q,z)=\Big{(}\tilde{\Psi}_{1}(\boldsymbol{a},h,q,z),\dots,\tilde{\Psi}_{n}(\boldsymbol{a},h,q,z)\Big{)},$
where
$\tilde{\Psi}_{i}(\boldsymbol{a},h,q,z)={\Psi}_{i}(\boldsymbol{a},h,q,z)\Phi(\boldsymbol{a},h,a_{i},q,z)$
and ${\Psi}_{i}(\boldsymbol{a},h,q,z)$ are given by (1.10). We can also rewrite the last formula simply as
$\tilde{\Psi}(\boldsymbol{a},h,q,z)={\Psi}(\boldsymbol{a},h,q,z)\Phi(\boldsymbol{a},h,x,q,z)$
where $\Phi(\boldsymbol{a},h,x,q,z)$ is now understood as the operator of multiplication by the gamma class (2.2) in the equivariant K - theory. In particular, in the basis of the $T$-fixed points this operator is represented by the diagonal matrix with eigenvalues the $\Phi(\boldsymbol{a},h,a_{i},q,z)$, $i=1,\dots,n$.
2.5.
The normalized fundamental solution matrices $\tilde{\Psi}(\boldsymbol{a},h,q,z)$ also have an enumerative meaning: these are the matrices of the capping operator for $X$, which is the partition function of the relative quasimaps to $X$, see Section 7.4 in [Oko17] for the definitions.
An important input from the enumerative geometry we need for our goals is the existence of the shift operators. The operators describe $q$-difference equations satisfied by the capping operator in the equivariant parameters $\boldsymbol{a}$ and $h$. These equations exist in general for an arbitrary Nakajima variety. Applied to $X=T^{*}\mathbb{P}^{n-1}$ this results can be formulated as follows:
Theorem 2.1 (Theorem 8.2.20, [Oko17]).
The normalized fundamental solution matrices satisfy the following system of difference equations
$\tilde{\Psi}(a_{1},\dots,a_{i}+1,\dots a_{n},h,q,z)=S^{a_{i}}(\boldsymbol{a},h,q,z)\tilde{\Psi}(a_{1},\dots,a_{i},\dots a_{n},h,q,z),$
$\tilde{\Psi}(a_{1},\dots,a_{i},\dots a_{n},h+1,q,z)=S^{h}(\boldsymbol{a},h,q,z)\tilde{\Psi}(a_{1},\dots,a_{i},\dots a_{n},h,q,z).$
where $S^{a_{i}}(\boldsymbol{a},h,q,z)$ and $S^{h}(\boldsymbol{a},h,q,z)$ denote matrices whose entries are rational functions in the equivariant parameters ${\boldsymbol{u}}$, $\hbar$, $q$ and $z$:
$S^{a_{i}}(\boldsymbol{a},h,q,z),S^{h}(\boldsymbol{a},h,q,z)\in GL_{n}(\mathbb{Q}({\boldsymbol{u}},\hbar,z,q)).$
3. $p$adic $q$-hypergeometric functions
3.1.
Let $p$ be a prime, $\mathbb{Q}_{p}$ be the filed of $p$-adic numbers, $\mathbb{Z}_{p}$ be the integers of $\mathbb{Q}_{p}$. Starting from this section $\mathbb{K}$ denotes the $p$-adic completion of the algebraic closure of $\mathbb{Q}_{p}$. Let us fix $q\in\mathbb{K}^{\times}$ of the following form
(3.1) $q=1+t,\ \ t\neq 0,\ \ |t|_{p}\leqslant 1.$
From now, we assume that all the equivariant parameters (1.14) are $p$ adic integers $a_{1},\dots,a_{n},h\in\mathbb{Z}_{p}$.
3.2.
Let $\overline{\mathbb{K}(z)}$ be the completion of $\mathbb{K}(z)$ with respect to the Gauss norm:
$\left|\sum\limits_{m=0}^{\infty}\,r_{m}(z)\right|_{\mathrm{Gauss}}=\ \ \max\,\{\,|r_{m}(z)|_{p}:m=0,1,2,\dots\}.$
A Frobenius intertwiner from $\mathcal{H}(\boldsymbol{a},h,q^{p},z^{p})$ to $\mathcal{H}(\boldsymbol{a}^{\prime},h^{\prime},q,z)$ is $\mathsf{U}(z)\in\mathsf{GL}_{n}(\overline{\mathbb{K}(z)})$ such that
$\mathbf{M}_{\mathcal{L}}(\boldsymbol{a}^{\prime},\hbar^{\prime},q,z)=\mathsf{U}(zq)^{-1}\,\mathbf{M}_{\mathcal{L}}(\boldsymbol{a},\hbar,q^{p},z^{p})\,\mathsf{U}(z).$
If ${\Psi}(\boldsymbol{a},\hbar,q^{p},z^{p})$ is a fundamental solution matrix for the hypergeometric system $\mathcal{H}(\boldsymbol{a},\hbar,q^{p},z^{p})$ (1.11) then $\mathsf{U}(z){\Psi}(\boldsymbol{a},\hbar,q^{p},z^{p})$ is a fundamental solution matrix for $\mathcal{H}(\boldsymbol{a}^{\prime},h^{\prime},q,z)$. Thus, it is of the form ${\Psi}(\boldsymbol{a}^{\prime},\hbar^{\prime},q,z)\Lambda$ and we obtain a relation
(3.2) ${\Psi}(\boldsymbol{a}^{\prime},\hbar^{\prime},q,z)\Lambda=\mathsf{U}(z){\Psi}(\boldsymbol{a},\hbar,q^{p},z^{p})$
for some constant matrix $\Lambda\in\mathsf{GL}_{n}(\mathsf{\mathbb{K}})$. The matrix $\mathsf{U}(0)$ is called the constant term of the Frobenius intertwiner. If a fundamental solution is normalized by ${\Psi}(\boldsymbol{a},\hbar,q,0)=1$, then $\mathsf{U}(0)=\Lambda$.
3.3.
For the purposes of this note, it is sufficient to restrict to a smaller ring $\widehat{\mathsf{R}}_{p}\subset\overline{\mathbb{K}(z)}$ which has a more explicit structure. Let $\mathsf{R}_{p}=\mathbb{Z}[z,(1-z)^{-1}]$ be the ring of rational functions with denominator given by powers of $(1-z)^{-1}$. The ring we need is the $p$-adic completion
$\widehat{\mathsf{R}}_{p}=\lim\limits_{\leftarrow}\,\mathsf{R}_{p}/p^{s}\mathsf{R}_{p}.$
By definition, the elements of $\widehat{\mathsf{R}}_{p}$ are sums of the form
$\sum\limits_{k=0}^{\infty}\,p^{k}\dfrac{g_{k}(z)}{(1-z)^{m_{k}}}$
where $m_{k}\in\mathbb{N}$ and $g_{k}(z)\in\mathbb{Z}[z]$. Put differently, an element $g(z)\in\widehat{\mathsf{R}}_{p}$ can be represented by a power series $g(z)\in\mathbb{Z}[[z]]$ such that for any $s\geqslant 0$ there is a natural number $m_{s}$ such that
$(1-z)^{m_{s}}g(z)\!\!\pmod{p^{s+1}}\in\mathbb{Z}[z].$
This condition provides a useful numerical test to determine whether a given power series $g(z)$ belongs to $\widehat{\mathsf{R}}_{p}$ or not, see Section 5. The Frobenius intertwiners we encounter in this paper are all $\mathsf{GL}_{n}(\widehat{\mathsf{R}}_{p})$ valued.
3.4.
For $n\in\mathbb{N}$ we define the $q$-number by:
$[n]_{q}=\dfrac{1-q^{n}}{1-q}=1+q+\dots+q^{n-1}.$
Let $[n]_{q}!=[1]_{q}[2]_{q}\dots[n]_{q}$ denote the $q$-factorial and
(3.3) $\binom{n}{k}_{\!\!q}=\dfrac{[n]_{q}!}{[k]_{q}![n-k]_{q}!}$
be the $q$-binomial factor. We recall the original definitions of the $p$-adic gamma and $q$-gamma functions:
Theorem 3.1 (Koblitz, [Ko80]).
For $q$ as in (3.1) the function $\Gamma_{p,q}(n)=(-1)^{n}\prod\limits_{{0<i<n}\atop{p\nmid i}}\,[n]_{q}$ extends to a continuous function on $\mathbb{Z}_{p}$. $\lim\limits_{q\to 1}\,\Gamma_{p,q}=\Gamma_{p}$ where $\Gamma_{p}$ denotes the $p$-adic gamma function of Morita [Mo75], which is the extension to $\mathbb{Z}_{p}$ of $\Gamma_{p}(n)=(-1)^{n}\prod\limits_{{0<i<n}\atop{p\nmid i}}\,n.$
3.5.
We formulate our main theorem in the geometric language of Section 2. For this, we replace the equivariant K-theory (2.1) by
(3.4) $K_{T}(X,\mathbb{K})=K_{T}(X)\otimes\mathbb{K}=\mathbb{K}[L^{\pm},u_{1}^{\pm},\dots,u^{\pm}_{n},\hbar^{\pm}]/(L-u_{1})\dots(L-u_{n}).$
The Frobenius intertwiner is understood as a map
$\mathsf{U}(z):K_{T}(X,\mathbb{K})[[z]]\longrightarrow K_{T}(X,\mathbb{K})[[z]],$
In particular, the constant term acts on $K$-theory
$\mathsf{U}(0):K_{T}(X,\mathbb{K})\longrightarrow K_{T}(X,\mathbb{K}).$
We are now ready to formulate our main result:
Theorem 3.2.
For any $\boldsymbol{a}=(a_{1},\dots,a_{n})\in\mathbb{Z}^{n}_{p},\ \ h\in\mathbb{Z}_{p}$ there exists unique Frobenius intertwiner $\mathsf{U}(z)\in\mathsf{GL}_{n}(\overline{\mathbb{K}(z)})$ from $\mathcal{H}(\boldsymbol{a},h,q^{p},z^{p})$ to $\mathcal{H}(p\boldsymbol{a},ph,q,z)$, where $p\boldsymbol{a}=(pa_{1},\dots,pa_{n})$. The constant term of the intertwiner acts in $K_{T}(X,\mathbb{K})$ as the operator of multiplication by the K-theory class
(3.5) $\mathsf{U}(0)=\prod\limits_{j=1}^{n}\,\dfrac{\Gamma_{p,q}(px+ph-pa_{j})}{\Gamma_{p,q}(px-pa_{j})\Gamma_{p,q}(ph)}.$
Clearly, multiplying a Frobenius intertwiner by a constant multiple $\mathsf{U}(z)\to c\mathsf{U}(z)$, $c\in\mathbb{K}^{\times}$ preserves all the properties. So, the “uniqueness” in Theorem 3.2 should be understood as uniqueness up to a scalar multiple.
If the torus fixed points of $T$-action are isolated, multiplication by any equivariant K-theory class is diagonal in the basis given by the classes of $T$-fixed points. In our case $X=T^{*}\mathbb{P}^{n-1}$ the $X^{T}$ is a finite set when all $a_{1},\dots,a_{n}$ are pairwise distinct. The corresponding fixed point component of $\mathsf{U}(0)$ are obtained by specialization $x=a_{i}$. We thus conclude with the following result.
Corollary 3.3.
If all $a_{1},\dots,a_{n}$ are pairwise distinct, the constant term of the Frobenius intertwiner $\mathsf{U}(0)$ acts in the basis of torus fixed points as a diagonal matrix with the following eigenvalues:
(3.6) $\mathsf{U}(0)_{i,i}=\prod\limits_{j=1}^{n}\,\dfrac{\Gamma_{p,q}(p(a_{i}-a_{j}+h))}{\Gamma_{p,q}(p(a_{i}-a_{j}))\Gamma_{p,q}(ph)}.$
Remark.
In the limit $q\to 1$ when $\Gamma_{p,q}\to\Gamma_{p}$ and Theorem 3.2 specializes to the results of Dwork [Dw89] for hypergeometric equations, see also [Ked21] for recent treatment. We consider this case in more details in Section 6.
4. Proof of Theorem 3.2
4.1.
Let $t\in\mathbb{Q}_{p}$ with $|t|_{p}\leqslant|p|_{p}=1/p$. Let $q=1+t$. Using elementary combinatorics one can derive the following congruence
(4.1) $(z,q)_{p^{s}}\equiv(z^{p},q^{p})_{p^{s-1}}\pmod{p^{s}}$
for instance, using the induction on $s$. Let us introduce the symbol
$[x,q]_{d}=(1-qx)\dots(1-q^{d-1}x)=\dfrac{(xq,q)_{\infty}}{(xq^{d},q)_{\infty}}$
which differs from the standard Pochhammer by a factor.
Lemma 4.1.
For any $i\in\mathbb{Z}$, modulo $p^{s}$ we have
$\dfrac{[q^{i},q]_{p^{s}}}{[1,q]_{p^{s}}}\equiv\left\{\begin{array}[]{ll}\dfrac{[q^{i},q^{p}]_{p^{s-1}}}{[1,q^{p}]_{p^{s-1}}},&\mathrm{if}\ \ p\mid i,\\ &\\ 0,&\mathrm{if}\ \ p\nmid i.\end{array}\right.$
Proof.
By the $q$-binomial theorem
$\sum\limits_{i=0}^{\infty}\,\dfrac{[q^{i},q]_{p^{s}}}{[1,q]_{p^{s}}}z^{i}=\dfrac{1}{(z,q)_{p^{s}}},\ \ \ \sum\limits_{i=0}^{\infty}\,\dfrac{[q^{pi},q^{p}]_{p^{s-1}}}{[1,q^{p}]_{p^{s-1}}}z^{pi}=\dfrac{1}{(z^{p},q^{p})_{p^{s-1}}},$
thus, for $i\geqslant$ the result follows from (4.1). Next for $-p^{s}+1\leqslant i\leqslant-1$
$[q^{i},q]_{p^{s}}=0$
and similarly for $[q^{i},q^{p}]_{p^{s-1}}$, thus, Lemma holds for these $i$’s too. Finally, if $i\leqslant-p^{s}$, the Lemma follows from the identity
$[q^{i},q]_{p^{s}}={q^{(p^{s}-1)i}q^{\frac{p^{s}(p^{s}-1)}{2}}}\,[q^{-(i+p^{s})},q]_{p^{s}}.$
4.2.
Let us consider the normalized vertex function (2.3) at $h=p^{s}$. We find
$\widetilde{V}(\boldsymbol{a},x,p^{s},q,z)=z^{x}\,\sum\limits_{d=0}^{\infty}\,\Big{(}\prod\limits_{i=1}^{n}\dfrac{[q^{\,x+d-a_{i}},q]_{p^{s}}}{[1,q]_{p^{s}}}\Big{)}z^{d}.$
Theorem 4.2.
We have the following congruences for the normalized K-theoretic vertex functions:
$\tilde{V}(p\boldsymbol{a},px,p^{s},q,z)\equiv\tilde{V}(\boldsymbol{a},x,p^{s-1},q^{p},z^{p})\pmod{p^{s}}.$
Proof.
The normalized K-theoretic vertex function is
$\tilde{V}(p\boldsymbol{a},px,p^{s},q,z)=z^{px}\,\sum\limits_{d=0}^{\infty}\,\left(\prod\limits_{i=1}^{n}\dfrac{[q^{px-pa_{i}+d},q]_{p^{s}}}{[1,q]_{p^{s}}}\right)z^{d}$
Lemma 4.1 gives
$z^{px}\,\sum\limits_{d=0}^{\infty}\,\left(\prod\limits_{i=1}^{n}\dfrac{[q^{px-pa_{i}+d},q]_{p^{s}}}{[1,q]_{p^{s}}}\right)z^{d}\equiv z^{px}\,\sum\limits_{d=0}^{\infty}\,\left(\prod\limits_{i=1}^{n}\dfrac{[q^{px-pa_{i}+pd},q^{p}]_{p^{s-1}}}{[1,q^{p}]_{p^{s-1}}}\right)z^{pd}\pmod{p^{s}}$
which is exactly the statement of the theorem. ∎
Since the vertex describes all solutions of the $q$-holonomic system we also obtain:
Corollary 4.3.
The normalized fundamental solution matrices satisfy the following congruences:
$\widetilde{\Psi}(p\boldsymbol{a},p^{s},q,z)\pmod{p^{s}}=\widetilde{\Psi}(\boldsymbol{a},p^{s-1},q^{p},z^{p}).$
4.3.
To prove Theorem 3.2 it is enough to show that for any $s\in\mathbb{N}$,
$\mathsf{U}(z)=\widetilde{\Psi}(p\boldsymbol{a},ph,q,z)\,\widetilde{\Psi}(\boldsymbol{a},h,q^{p},z^{p})^{-1}\pmod{p^{s}},$
is a rational function in $z$. Let us write $h=h_{0}+p^{s-1}h_{1}$ for some $h_{0}\in\mathbb{N}$ and $h_{i}\in\mathbb{Z}_{p}$. By Theorem 2.1 we have:
$\mathsf{U}(z)=\widetilde{\Psi}(p\boldsymbol{a},ph,q,z)=\widetilde{\Psi}(p\boldsymbol{a},ph_{0}+h_{1}p^{s},q,z)=\left(\prod\limits_{i=h_{1}p^{s}}^{ph_{0}+h_{1}p^{s}-1}S^{h}(p\boldsymbol{a},i,q,z)\right)\widetilde{\Psi}(p\boldsymbol{a},h_{1}p^{s},q,z),$
where $S^{h}(p\boldsymbol{a},i,q,z)$ is a certain invertible matrix whose coefficients are rational functions of $z$. By Corollary 4.3 we obtain:
$\widetilde{\Psi}(p\boldsymbol{a},ph,q,z)\equiv\left(\prod\limits_{i=0}^{ph_{0}-1}S^{h}(p\boldsymbol{a},i,q,z)\right)\widetilde{\Psi}(\boldsymbol{a},h_{1}p^{s-1},q^{p},z^{p})\pmod{p^{s}}.$
Similarly, we obtain
$\widetilde{\Psi}(\boldsymbol{a},h,q^{p},z^{p})\equiv\left(\prod\limits_{i=0}^{h_{0}-1}S^{h}(\boldsymbol{a},i,q^{p},z^{p})\right)\widetilde{\Psi}(\boldsymbol{a},h_{1}p^{s-1},q^{p},z^{p})$
and thus
(4.2) $\widetilde{\Psi}(p\boldsymbol{a},ph,q,z)\,\widetilde{\Psi}(\boldsymbol{a},h,q^{p},z^{p})^{-1}\equiv\left(\prod\limits_{i=0}^{ph_{0}-1}S^{h}(p\boldsymbol{a},i,q,z)\right)\left(\prod\limits_{i=0}^{h_{0}-1}S^{h}(\boldsymbol{a},i,q^{p},z^{p})\right)^{-1}\pmod{p^{s}}$
which is a product of finitely many matrices whose entries are rational functions in $z$.
4.4.
Finally, from the normalization (2.2), (2.3) we find:
$\mathsf{U}(0)=\dfrac{\Phi(p\boldsymbol{a},ph,px,q,z)}{\Phi(\boldsymbol{a},h,x,q^{p},z^{p})}$
From definition (2.2) we see that $\mathsf{U}(0)$ is exactly the product of factors as in Lemma 4.4 below, and therefore
$\mathsf{U}(0)=\prod\limits_{j=1}^{n}\,\dfrac{\Gamma_{p,q}(px+ph-pa_{j})}{\Gamma_{p,q}(px-pa_{j})\Gamma_{p,q}(ph)}$
which finishes the proof of the theorem.
4.5.
If $n$ and $h$ are two natural numbers then
$\dfrac{(q^{ph},q)_{\infty}}{(q,q)_{\infty}}\dfrac{(q^{pn+1},q)_{\infty}}{(q^{pn+ph},q)_{\infty}}=\binom{pn+ph-1}{pn}_{\!\!q}.$
is a $q$-binomial coefficient (3.3). Similarly
$\dfrac{(q^{ph},q^{p})_{\infty}}{(q^{p},q^{p})_{\infty}}\dfrac{(q^{pn+p},q^{p})_{\infty}}{(q^{pn+ph},q^{p})_{\infty}}=\binom{n+h-1}{n}_{\!\!q^{p}}.$
For the ratio of these factors we find:
$\dfrac{\binom{pn+ph-1}{pn}_{\!\!q}}{\binom{n+h-1}{n}_{\!\!q^{p}}}=\dfrac{\prod\limits_{{0<i<pn+ph}\atop{p\nmid i}}\,[i]_{q}}{(\prod\limits_{{0<i<pn}\atop{p\nmid i}}\,[i]_{q})(\prod\limits_{{0<i<ph}\atop{p\nmid i}}\,[i]_{q})}=\dfrac{\Gamma_{p,q}(pn+ph)}{\Gamma_{p,q}(pn)\,\Gamma_{p,q}(ph)}$
where the last equality is by definition of $(p,q)$-Gamma function. By Koblitz’s Theorem 3.1 we obtain
Lemma 4.4.
The function of $n,h\in\mathbb{N}$ given by
$\dfrac{\dfrac{(q^{ph},q)_{\infty}}{(q,q)_{\infty}}\dfrac{(q^{pn+1},q)_{\infty}}{(q^{pn+ph},q)_{\infty}}}{\dfrac{(q^{ph},q^{p})_{\infty}}{(q^{p},q^{p})_{\infty}}\dfrac{(q^{pn+p},q^{p})_{\infty}}{(q^{pn+ph},q^{p})_{\infty}}}=\dfrac{\Gamma_{p,q}(pn+ph)}{\Gamma_{p,q}(pn)\,\Gamma_{p,q}(ph)}$
extends to $\mathbb{Z}_{p}$ values of arguments $n$ and $h$.
5. Numerical examples of Theorem 3.2
5.1.
Let ${\Psi}(\boldsymbol{a},\hbar,q,z)$ be the fundamental solution matrix (1.12) for some fixed parameters. As in (3.2) we look for the Frobenius intertwiner matrix in the form
(5.1) $\mathsf{U}(z)={\Psi}(p\boldsymbol{a},ph,q,z)\Lambda{\Psi}(\boldsymbol{a},h,q^{p},z^{p})^{-1}.$
where $\Lambda$ is a constant matrix. In order to be a Frobenius intertwiner, we need to find such $\Lambda\in\mathsf{GL}_{n}(\mathbb{K})$ for which (5.1) is an element of $\mathsf{GL}_{n}(\widehat{\mathsf{R}}_{p})$. Explicitly, $\mathsf{U}(z)$ must be a power series
$\mathsf{U}(z)=\mathsf{U}_{0}+\mathsf{U}_{1}z+\dots\in\mathsf{Mat}_{n}(\mathbb{Z})[[z]]$
such that for any natural number $s$, there exists a natural number $m_{s}$ for which
(5.2) $(1-z)^{m_{s}}\,\mathsf{U}(z)\pmod{p^{s}}\in\mathsf{Mat}_{n}(\mathbb{Z})[z].$
In words: after multiplication by sufficiently large power of $(1-z)$ the matrix elements of $\mathsf{U}(z)$ are polynomials in $z$ modulo $p^{s}$. This condition is very strong and can be used to effectively compute the Frobenius intertwiners as we about to show.
5.2.
Note that if the parameters $\boldsymbol{a}=(a_{1},\dots,a_{n})$ are pairwise distinct the matrix $\Lambda$ must be diagonal. Indeed, the fundamental solution matrix $\tilde{\Psi}(\boldsymbol{a},\hbar,q,z)$ contains logarithmic terms given by matrix $\mathsf{E}(z)$ in (1.13). The intertwiner (5.1) is a power series in $z$ only if these log terms are cancelled. It is elementary to check that for the pairwise distinct $a_{i}$ this happens only if $\Lambda$ is diagonal. Since $\Lambda$ is defined up to a scalar factor, we may assume that $\Lambda_{n,n}=1$ so
(5.8) $\Lambda=\left(\begin{array}[]{ccccc}c_{1}&0&0&..&0\\ 0&c_{2}&0&..&0\\ ..&..&..&..&..\\ 0&0&..&c_{n-1}&0\\ 0&0&0&..&1\end{array}\right)$
for some unknown $c_{1},\dots,c_{n-1}\in\mathbb{Z}_{p}$.
5.3.
Let us assume $n=2$. In this case
(5.11) $\Lambda=\left(\begin{array}[]{cc}c&0\\ 0&1\end{array}\right).$
Let us pick a prime $p=3$. To meet the condition (3.1) let us also set
$q=1+p=4,$
and pick some random values of $a_{1},a_{2},h\in\mathbb{Z}_{p}$, for instance
(5.12) $h=\dfrac{1}{2},\ \ a_{1}=\dfrac{1}{5},\ \ a_{2}=0.$
(Since the solutions only depend on the difference $a_{1}-a_{2}$ we can always assume $a_{2}=0$). For our computer check we use integral approximations of these numbers say up to $p^{7}$:
$\bar{h}=h\pmod{p^{7}}=1094,\ \ \bar{a_{1}}=a_{1}\pmod{p^{7}}=875,\ \ \ \bar{a_{2}}=0.$
We now determine the constant $c\in\mathbb{Z}_{p}$ in (5.11) from the rationality condition (5.2) in a chain of steps. For this we assume that $c$ has the following $p$-adic expansion:
$c=c^{(0)}+c^{(1)}3+c^{(2)}3^{2}+\dots$
where $c^{(i)}\in\{0,1,2\}$.
Step 1: determination of $c^{(0)}$ and $c^{(1)}.$
Using a computer we approximate the power series ${\Psi}(\boldsymbol{a},\hbar,q,z)$ by expansion up to $O(z^{20})$. Substituting this approximation to (5.1) and trying the integers $0,1,2$ for the values of $c^{(0)}$ and $c^{(1)}$ we find that
$\mathsf{U}(z)\pmod{p}$
exists (i.e. no powers of $p$ in the denominators) only if $c^{(0)}=1$ and $c^{(1)}=0$. In which case we find that
(5.15) $\mathsf{U}(z)\pmod{p}=\left(\begin{array}[]{cc}1+O\left({z}^{20}\right)&O\left({z}^{20}\right)\\ \par\vskip 6.0pt plus 2.0pt minus 2.0pt\cr O\left({z}^{20}\right)&1+O\left({z}^{20}\right)\end{array}\right)$
We note that $\mathsf{U}(z)\pmod{p}$ given by (5.15) does not depend on the value of $c^{(2)}$ and higher.
Step 2: determination of $c^{(2)}.$
If $c^{(2)}=0$, using a computer we find
$\mathsf{U}(z)\pmod{p}=\left[\begin{array}[]{cc}1+O\left({z}^{20}\right)&O\left({z}^{20}\right)\\ \par\vskip 6.0pt plus 2.0pt minus 2.0pt\cr{z}^{12}+2\,{z}^{15}+O\left({z}^{20}\right)&1+O\left({z}^{20}\right)\end{array}\right]$
If $c^{(2)}=1$ we find
$\mathsf{U}(z)\pmod{p}=\left[\begin{array}[]{cc}1+O\left({z}^{20}\right)&O\left({z}^{20}\right)\\ \par\vskip 6.0pt plus 2.0pt minus 2.0pt\cr O\left({z}^{20}\right)&1+O\left({z}^{20}\right)\end{array}\right]$
If $c^{(2)}=2$ we find
$\mathsf{U}(z)\pmod{p}=\left[\begin{array}[]{cc}1+O\left({z}^{20}\right)&O\left({z}^{20}\right)\\ \par\vskip 6.0pt plus 2.0pt minus 2.0pt\cr 2\,{z}^{12}+{z}^{15}+O\left({z}^{20}\right)&1+O\left({z}^{20}\right)\end{array}\right]$
We see that the only value of $c^{(2)}$ for which the result is a polynomial in $z$ is $c^{(2)}=1$. These results do not depend on the value of the coefficient $c^{(3)}$ and higher which are to be determined in the next step.
Step 3: determination of $c^{(3)}.$
If $c^{(3)}=0$ we find:
$(1-z)^{2}\mathsf{U}(z)\!\!\pmod{p^{2}}=\left[\begin{array}[]{cc}1+z+{z}^{2}+O\left({z}^{20}\right)&O\left({z}^{20}\right)\\ \par\vskip 6.0pt plus 2.0pt minus 2.0pt\cr 6\,z+3\,{z}^{12}+3\,{z}^{13}+3\,{z}^{14}+6\,{z}^{15}+6\,{z}^{16}+6\,{z}^{17}+O\left({z}^{20}\right)&1+7\,z+{z}^{2}+O\left({z}^{20}\right)\end{array}\right]$
If $c^{(3)}=1$ we find:
$(1-z)^{2}\mathsf{U}(z)\!\!\pmod{p^{2}}=\left[\begin{array}[]{cc}1+z+{z}^{2}+O\left({z}^{20}\right)&O\left({z}^{20}\right)\\ \par\vskip 6.0pt plus 2.0pt minus 2.0pt\cr 6\,z+O\left({z}^{20}\right)&1+7\,z+{z}^{2}+O\left({z}^{20}\right)\end{array}\right]$
If $c^{(3)}=2$ we find:
$(1-z)^{2}\mathsf{U}(z)\!\!\pmod{p^{2}}=\left[\begin{array}[]{cc}1+z+{z}^{2}+O\left({z}^{20}\right)&O\left({z}^{20}\right)\\ \par\vskip 6.0pt plus 2.0pt minus 2.0pt\cr 6\,z+6\,{z}^{12}+6\,{z}^{13}+6\,{z}^{14}+3\,{z}^{15}+3\,{z}^{16}+3\,{z}^{17}+O\left({z}^{20}\right)&1+7\,z+{z}^{2}+O\left({z}^{20}\right)\end{array}\right]$
We conclude that there is unique value of $c^{(3)}=1$ for which the result is a polynomial in $z$.
Continuing this process we will find that at $m$-th step there is always unique value of $c^{m}$ for which rationality condition (5.2) is met. For instance, up to the order $9$ these method gives the following value
[…]
5.4.
Theorem 3.2 and Corollary 3.3 gives the following formula for the constant $c$:
$c=\dfrac{\Lambda_{1,1}}{\Lambda_{2,2}}=\dfrac{\Gamma_{p,q}(p(a_{1}-a_{2}+h))\Gamma_{p,q}(p(a_{2}-a_{1}))}{\Gamma_{p,q}(p(a_{2}-a_{1}+h))\Gamma_{p,q}(p(a_{2}-a_{1}))}$
at (5.12) we have
$c=\dfrac{\Gamma_{p,q}(\frac{21}{10})\Gamma_{p,q}(-\frac{3}{5})}{\Gamma_{p,q}(\frac{9}{10})\Gamma_{p,q}(\frac{3}{5})}.$
For $p=3$ and $q=1+p=4$ we compute:
[…]
in full agreement with the value obtained from the numerical experiment (5.16).
The author have performed the same numerical experiments for numerous other rational values of parameters $a_{1},a_{2}$ and $h$ in $\mathbb{Z}_{p}$. In all these cases the results are in agreement with Theorem 3.2. Same method can be used for ranks $n>2$, in which case we have to determine the numerical values of $n-1$ constants in (5.8). Again, we find full agreement with Theorem 3.2.
6. Limit to cohomology
6.1.
As $q\to 1$, the quantum difference equation of $X$ degenerates to the quantum differential equation, which governs quantum cohomology of $X$. In this limit, the equivairnat K-theory ring (2.1) is replaced by the equivairnat cohomology
(6.1) $H^{\bullet}_{T}(X)=\mathbb{Z}[x,a_{1},\dots,a_{n},h]/(x-a_{1})\dots(x-a_{n}).$
In this limit the gamma class (2.2) takes the form:
(6.2) $\Gamma(\boldsymbol{a},h,x,z)=z^{x}\,\prod\limits_{i=1}^{n}\dfrac{\Gamma(x+h-a_{i})}{\Gamma(x+1-a_{i})\Gamma(h)}\in\hat{H}^{\bullet}_{T}(X).$
where $\Gamma$ denotes the classical Gamma function, and $\hat{H}^{\bullet}_{T}(X)$ is the appropriate completion of the cohomology ring containing the gamma class. The normalized cohomological vertex function (also known as $J$-function in cohomology) is:
$\widetilde{V}(\boldsymbol{a},h,x,z)=\sum\limits_{d=0}^{\infty}\,\Gamma(\boldsymbol{a},h,x+d,z)\,\in\,\hat{H}^{\bullet}_{T}(X)[[z]].$
The vertex function satisfy the differential equation
(6.3) $P(\boldsymbol{a},h,z)\widetilde{V}(\boldsymbol{a},h,x,z)=(x-a_{1})\dots(x-a_{n})\Gamma(\boldsymbol{a},h,x,z)=0$
(note that the right side is zero in (6.1)), where
(6.4) $P(\boldsymbol{a},h,z)=z\prod\limits_{i=1}^{n}\,(D+h-a_{i})-\prod\limits_{i=1}^{n}(D-a_{i})$
and $D=z\frac{d}{dz}$. The components of the vertex in some basis of $H^{\bullet}_{T}(X)$ provide a basis of solutions for the hypergeometric differential equation (6.3). In particular, in the basis given by the classes of the $T$-fixed points $X^{T}$, we obtain the standard generazlized hypergeometric series:
$\widetilde{V}(\boldsymbol{a},h,a_{i},z)=\Gamma(\boldsymbol{a},h,a_{i},z)\,\sum\limits_{d=0}^{\infty}\,\left(\prod\limits_{j=1}^{n}\dfrac{(a_{i}+h-a_{j})_{d}}{(a_{i}+1-a_{j})_{d}}\right)z^{d}$
where $(a)_{d}=a(a+1)\dots(a+d-1)$.
By Theorem 3.1 $\Gamma_{p,q}(n)\stackrel{{\scriptstyle q\to 1}}{{\longrightarrow}}\Gamma_{p}(n)$ and from Theorem 3.2 we deduce:
Theorem 6.1.
For any $\boldsymbol{a}=(a_{1},\dots,a_{n})\in\mathbb{Z}^{n}_{p},\ \ h\in\mathbb{Z}_{p}$ there exists unique Frobenius intertwiner $\mathsf{U}(z)$ from the hypergeometric system (6.4) with parameters $(a_{1},\dots,a_{n},h,z^{p})$ to the system with parameters $(pa_{1},\dots,pa_{n},ph,z)$. The constant terms of the intertwiner is represented by the operator acting in $\hat{H}^{\bullet}_{T}(X)$ as an operator of multiplication by the cohomology class class
(6.5) $\mathsf{U}(0)=\prod\limits_{j=1}^{n}\,\dfrac{\Gamma_{p}(px+ph-pa_{j})}{\Gamma_{p}(px-pa_{j})\Gamma_{p}(ph)},$
where $\Gamma_{p}$ is the Morita gamma function.
The analog of Corollary 3.3 also gives:
Corollary 6.2.
If $a_{1},\dots,a_{n}$ are pairwise distinct, the constant term of the Frobenius intertwiner in the basis of torus fixed points is given by diagonal matrix with the following eigenvalues
(6.6) $\mathsf{U}(0)_{i,i}=\prod\limits_{j=1}^{n}\,\dfrac{\Gamma_{p}(p(a_{i}-a_{j}+h))}{\Gamma_{p}(p(a_{i}-a_{j}))\Gamma_{p}(ph)}.$
In this way we arrive at the description of the Frobenius intertwines for the $p$-adic hypergeometric equations discovered by Dwork [Dw89] see also [Ked21] where this result was recently revisited by other tools.
Remark.
Recall that for $n\in p\mathbb{Z}_{p}$ the Morita Gamma function satisfies the identities $\Gamma_{p}(n)=-\Gamma_{p}(1+n)=(-1)^{n_{0}}/\Gamma_{p}(-n)$. Using this, up to a scalar multiple, the constant term of Frobenius (6.5) takes the form
$\mathsf{U}(0)=\prod\limits_{i=1}^{n}\,\Gamma_{p}(px+ph-pa_{j})\Gamma_{p}(pa_{j}-px)$
The arguments of gamma functions in this product are exactly the $p$-multiples of the Chern roots of the tangent bundle $TX$. We thus can write this formally as
(6.7) $\mathsf{U}(0)=\Gamma_{p}(TX^{(p)}),$
where $(p)$ indicates the Frobenius twist multiplying the Chern roots by $p$. We expect that (6.7) holds for the constant term of the Frobenius intertwiner of the quantum differential equation for an arbitrary $X$.
7. Limit to $X=\mathbb{P}^{n-1}$: Frobenius structure for generalized Bessel equations
7.1.
The quantum differential equation and its solutions for $X=\mathbb{P}^{n-1}$ can be obtained as the scaling limit $\hbar\to\infty$ of the corresponding equations and solutions for $T^{*}\mathbb{P}^{n-1}$ (we recall that the equivariant parameter $\hbar$ is the character corresponding to the dilation of the cotangent directions of $T^{*}\mathbb{P}^{n-1}$ under the torus action) [SV24]. In this section we consider the Frobenius intertwiners for the quantum differential equation of $\mathbb{P}^{n-1}$. We also describe the Frobenius automorphism for the non-equivariant limit of this equations which correspond to the generalized Bessel equations.
7.2.
As before, let $T=(\mathbb{C}^{\times})^{n}$ be the torus acting on $\mathbb{C}^{n}$ by scaling the coordinate lines with characters $a_{1},\dots,a_{n}$. We consider the induced action on $X=\mathbb{P}^{n-1}=\mathbb{P}(\mathbb{C}^{n})$. The equivariant cohomology ring has the form
(7.1) $H^{\bullet}_{T}(X)=\mathbb{Z}[x,a_{1},\dots,a_{n}]/(x-a_{1})\dots(x-a_{n}).$
Taking the formal limit $\hbar\to\infty$ of (6.2) we obtain the gamma class for $X=\mathbb{P}^{n-1}$:
$\Gamma(x,\boldsymbol{a},z)=\dfrac{z^{x}}{\prod\limits_{i=1}^{n}\Gamma(x+1-a_{i})}\in\hat{H}^{\bullet}_{T}(X).$
The normalized equivariant cohomological vertex function (the $J$-function) of $\mathbb{P}^{n-1}$ equals:
$\widetilde{V}(x,\boldsymbol{a},z)=\sum\limits_{d=0}^{\infty}\,\Gamma(x+d,\boldsymbol{a},z)\,\in\,\hat{H}^{\bullet}_{T}(X)[[z]].$
More explicitly
$\widetilde{V}(x,\boldsymbol{a},z)=\dfrac{z^{x}}{\prod\limits_{i=1}^{n}\Gamma(x+1-a_{i})}\,\sum\limits_{d=0}^{\infty}\,\dfrac{z^{d}}{(x-a_{1}+1)_{d}\dots(x-a_{n}+1)_{d}}.$
This power series satisfies the differential equation
$P(\boldsymbol{a},z)\widetilde{V}(x,\boldsymbol{a},z)=(x-a_{1})\dots(x-a_{n})\Gamma(x,\boldsymbol{a},z),$
where $P(\boldsymbol{a},z)$ is the differential operator
(7.2) $P(\boldsymbol{a},z)=z-(D-a_{1})\cdots(D-a_{n}),\ \ \ D=z\dfrac{d}{dz}$
and therefore in the ring (7.1) we have
$P(\boldsymbol{a},z)\widetilde{V}(x,\boldsymbol{a},z)=0.$
Formal limit $h\to\infty$ of Theorem 6.1 then gives:
Theorem 7.1.
[…]
(7.3) $\mathsf{U}(0)=\prod\limits_{j=1}^{n}{\Gamma_{p}}(px-pa_{j}).$
We note that the arguments of the gamma functions in the product (7.3) are the $p$ multiples of the Chern roots of the tangent bundle $TX$. Thus, the universal formula (6.7) holds, in agreement with the Remark of the previous Section.
7.3.
At $\boldsymbol{a}=(0,\dots,0)$ the equivariant cohomology of $X=\mathbb{P}^{n-1}$ specializes at the usual (non-equivariant) cohomology:
(7.4) $H^{\bullet}(X)=\mathbb{Z}[x]/(x)^{n}.$
The gamma class and the vertex function have the form
$\Gamma(x,z)=\dfrac{z^{x}}{\Gamma(x+1)^{n}},$
and
(7.5) $\widetilde{V}(x,z)=\sum\limits_{d=0}^{\infty}\,\Gamma(x+d,z)=\dfrac{z^{x}}{\Gamma(x+1)^{n}}\,\sum\limits_{d=0}^{\infty}\,\dfrac{z^{d}}{(x+1)^{n}_{d}},$
respectively. The vertex function solves
$P(z)\widetilde{V}(x,z)=x^{n}\,\Gamma(x,z)=0$
where
(7.6) $P(z)=z-D^{n}$
The differential equation defined by (7.6) is the generalized Bessel equation. For $n=2$ this is the classical Bessel equation. The components of $\widetilde{V}(x,z)$ in any basis of cohomology (7.4) provide a basis of solutions for these equations. In this case, the only natural basis is given by the powers of the hyperplane class $x$:
(7.7) Basis $\textrm{Basis of}\ \ H^{\bullet}(X)=\{1,x,x^{2},\dots,x^{n-1}\}.$
In order to compute the corresponding components it is enough to expand (7.5) in the Taylor expansion up to degree $x^{n}$:
$\widetilde{V}(x,z)=\widetilde{V}_{0}(z)+\widetilde{V}_{1}(z)x+\dots+\widetilde{V}_{n-1}x^{n-1}$
The functions $\widetilde{V}_{0}(z),\dots,\widetilde{V}_{n}(z)$ provide a basis of solutions to the generalized Bessel equation. Note that (7.5) has a multiple $z^{x}$ which has the following expansion:
$z^{x}=\sum\limits_{m=0}^{n-1}\,\dfrac{\ln(z)^{m}}{m!}x^{m}$
and thus the solutions may be written in the form
(7.13) $(\widetilde{V}_{0}(z),\dots,\widetilde{V}_{n-1}(z))=({V}_{0}(z),\dots,{V}_{n-1}(z))\cdot\left(\begin{array}[]{ccccc}1&\ln(z)&\ln(z)^{2}/2&..&\ln(z)^{n-1}/(n-1)!\\ 0&1&\ln(z)&..&\ln(z)^{n-2}/(n-2)!\\ 0&0&1&..&\dots\\ ..&..&..&..&\dots\\ 0&0&0&..&1\end{array}\right)$
where $V_{i}(z)$ are some analytic (power series in $z$) functions. In particular, $\tilde{V}_{0}(z)$ does not contain logarithmic terms. For $n=2$, $\tilde{V}_{0}(z)$ is the power series representing the classical Bessel function of the first kind.
7.4.
When all $a_{i}=0$, the Theorem 7.3 specializes to:
Theorem 7.2.
The generalized Bessel equation admits Frobenius intertwiner $z\to z^{p}$ with the constant term acting in $H^{\bullet}(X)$ as the operator of multiplication by the cohomology class
(7.14) $\mathsf{U}(0)={\Gamma_{p}}(px)^{n}$
By construction, the intertwiner of Theorem 7.3 maps the solutions $\tilde{V}(x,z^{p})$ to the solutions $\tilde{V}(px,z)$. Let $p^{\mathrm{deg}}$ be the operator of the cohomological grading, acting on the cohomology by $p^{\mathrm{deg}}(x^{k})=p^{k}x^{k}$. The above theorem says that
${\Gamma_{p}}(x)^{n}\cdot p^{\mathrm{deg}}$
is the zero term of the intertwiner between $\tilde{V}(x,z^{p})$ and $\tilde{V}(x,z)$ (note that as in (7.13), the operators act by multiplication from the right in our conventions). In other words this gives an automorphism of the corresponding differential equation. We thus conclude with
Theorem 7.3.
The generalized Bessel equation admits unique Frobenius automorphism acting by $z\to z^{p}$. The constant term of the automorphsim acts in cohomology as the operator
(7.15) $\mathsf{U}(0)={\Gamma_{p}}(x)^{n}\cdot p^{\mathrm{deg}}.$
Let us compute the matrix (7.15) explicitly in the basis (7.7). If $\Psi(x,z)=\Psi_{0}(z)+\Psi_{1}(z)x+\dots+\Psi_{n-1}x^{n-1}$, then the convention is that the matrices of the operators act on the row vector $(\Psi_{0}(z),\Psi_{1}(z),\dots,\Psi_{n-1})$ by multiplication from the right. For instance, multiplication by $x$ acts by $x\Psi(x,z)=\Psi_{0}(z)x+\Psi_{1}(z)x^{2}+\dots+\Psi_{n-2}x^{n-1}$, and thus the corresponding matrix is
$x=\left(\begin{array}[]{cccccc}0&1&0&0&\dots&0\\ 0&0&1&0&\dots&0\\ 0&0&0&1&\dots&0\\ ..&..&..&..&\dots&1\\ 0&0&0&0&\dots&0\\ \end{array}\right).$
Let us recall that the logarithm of the Morita gamma function has the following Taylor expansion, see for instance Sections 58 and 61 in [Sh]:
$\log\Gamma_{p}(x)=\Gamma_{p}^{\prime}(0)x-\sum\limits_{m\geqslant 2}\,\dfrac{\zeta_{p}(m)}{m}\,x^{m}$
where $\zeta_{p}(m)$ denote the values of $p$-adic zeta function. It is known that $\zeta_{p}(m)=0$ for even $m$, so only the odd powers of $x$ appear in the expansion. Thus, in our conventions we obtain that the operator of multiplication by $\log\Gamma_{p}(x)^{n}$ is given by the $n\times n$ matrix
$\log\Gamma_{p}(x)^{n}=\left(\begin{array}[]{cccccc}0&\nu_{1}&0&\nu_{3}&\dots&\dots\\ 0&0&\nu_{1}&0&\nu_{3}&\dots\\ 0&0&0&\nu_{1}&\dots&\dots\\ ..&..&..&..&\dots&0\\ ..&..&..&..&\dots&\nu_{1}\\ 0&0&0&0&\dots&0\\ \end{array}\right),$
where we denoted $\nu_{1}=n\Gamma_{p}^{\prime}(0)$ and $\nu_{m}=-n\,\zeta_{p}(m)/m$ for $m>2$. We also obviously have
$p^{\mathrm{deg}}=\left(\begin{array}[]{cccccc}1&0&0&0&\dots&0\\ 0&p&0&0&\dots&0\\ 0&0&p^{2}&0&\dots&0\\ ..&..&..&..&\dots&0\\ ..&..&..&..&p^{n-2}&0\\ 0&0&0&0&\dots&p^{n-1}\\ \end{array}\right).$
Finally, we obtain that the constant term of the automorphism has the following matrix:
$\mathsf{U}(0)={\Gamma_{p}}(x)^{n}\cdot p^{\mathrm{deg}}=\exp\left(\begin{array}[]{cccccc}0&\nu_{1}&0&\nu_{3}&\dots&\dots\\ 0&0&\nu_{1}&0&\nu_{3}&\dots\\ 0&0&0&\nu_{1}&\dots&\dots\\ ..&..&..&..&\dots&0\\ ..&..&..&..&\dots&\nu_{1}\\ 0&0&0&0&\dots&0\\ \end{array}\right)\cdot\left(\begin{array}[]{cccccc}1&0&0&0&\dots&0\\ 0&p&0&0&\dots&0\\ 0&0&p^{2}&0&\dots&0\\ ..&..&..&..&\dots&0\\ ..&..&..&..&p^{n-2}&0\\ 0&0&0&0&\dots&p^{n-1}\\ \end{array}\right)$
and the exponent of a nilpotent operator is, of course, elementary to compute.
Example.
For the case of the classical Bessel equation, corresponding to $n=2$, this matrix has the form
(7.18) $\mathsf{U}(0)=\left(\begin{array}[]{ll}1,&p\,\Gamma^{\prime}_{p}(0)\\ 0,&p\end{array}\right).$
The existence of the Frobenius automorphism for the Bessel equation was discovered by B. Dwork [Dwo74]. In particular, in Lemma 3.2 of [Dwo74] he gave an explicit formula for the matrix $\mathsf{U}(0)$ in terms of certain converging $p$-adic series. One checks that Dwork’s matrix $\mathfrak{U}(0)$ coincides with (7.18) after transposition. In particular, the series for the coefficients $\mathfrak{U}_{2,1}(0)=\gamma$ converges to $p\Gamma^{\prime}_{p}(0)$.
The action of the Frobenius automorphism for the generalized Bessel equations ($n>2$) was further studied by S. Sperber in [Sp80]. The results of this Section are in full agreement with these calculations too.
8. Frobenius intertwiners at $p^{s}$ roots of unity
8.1.
Let $\mathsf{U}(z)$ be the Frobenius intertwiner for the $q$-difference equation as in Theorem 3.2. As we expain in Section 3.2, the defining property of the Frobenius intertwiner $\mathsf{U}(z)$ discussed in the previous Sections is the condition
$\mathsf{U}(z)\pmod{p^{s}}={\mathrm{rational\ function\ of}}\ \ z$
In fact, a stronger property holds
$\mathsf{U}(z)\pmod{[p^{s}]_{q}}={\mathrm{rational\ function\ of}}\ \ z,$
where
$[p^{s}]_{q}=\dfrac{1-q^{p^{s}}}{1-q}=\varphi_{p}(q)\varphi_{p}(q^{p})\dots\varphi_{p}(q^{p^{s-1}})$
and $\varphi_{p}(q)=1+q+\dots+q^{p-1}$ is the $p$ th cyclotomic polynomial. This result follows from:
Theorem 8.1.
For $s>0$ let $\zeta\in\mathbb{K}$ denotes a primitive $p^{s}$ th roots of unity, then
$\left.\mathsf{U}(z)\right|_{q=\zeta}$
is well defined and is a rational function of $z$.
Proof.
At $q=\zeta$ we have the following (obvious) identity
$(z,q)_{p^{s}}=(z^{p},q^{p})_{p^{s-1}}=(1-z^{p^{s}})$
Arguing as in the proof of Lemma 4.1 we would find that at $q=\zeta$:
$\dfrac{[q^{i},q]_{p^{s}}}{[1,q]_{p^{s}}}=\left\{\begin{array}[]{ll}\dfrac{[q^{i},q^{p}]_{p^{s-1}}}{[1,q^{p}]_{p^{s-1}}}=1,&\mathrm{if}\ \ p\mid i,\\ &\\ 0,&\mathrm{if}\ \ p\nmid i.\end{array}\right.$
We then use this result as the analog of Lemma 4.1 to prove that the intertwiner (4.2) is rational, following the same steps as in the proof of Theorem 3.2. ∎
Remark.
For the case $n=1$ this result was recently obtained by D. Vargas-Montoya in [VM22]. In the terminology of [VM22] Theorem 8.1 can be formulated as existence of the strong Frobenius structure for the $q$-hypergeometric equation (1.2).
9. Concluding remarks and future directions
9.1.
It is known that the Nakajima quiver varieties have very rich enumerative geometry [Oko17]. Partition functions of quasimaps to the Nakajima quiver varieties with various boundary conditions satisfy remarkable systems of $q$-difference equations, known as the quantum difference equations [OS22]. In the present paper we discuss the existence of the Frobenuis structure for the quantum difference equations associated with simplest Nakajima variety $X=T^{*}\mathbb{P}^{n-1}$. All essential ingredients in our construction, e.g., the existence of the rational shift operators in the equivariant parameter $\hbar$, exist for general Nakajima varieties too. For this reason, the $p$ adic quantum difference equations of the Nakajima varieties are equipped with the Frobenius structure. We plan to return to the case of general $X$ in a separate note. In the cohomological limit $q\to 1$ the quantum difference equations degenerate to the quantum differential equations for the Nakajima varieties [MO19], therefore the Frobenius structures exists for such differential equations also.
9.2.
The quantum difference/differential equations of the Nakajima varieties are connected deeply with representation theory of quantum groups. It is an interesting problem to find a representation theoretic description of the Frobenius intertwiners for these equations. In particular, the quantum difference equations for $q$ specialised at the roots of unity are controlled by the Bethe ansatz, see [AO17, PSZ, KPSZ]. It seems possible to describe the Frobenius intertwiners for these values of $q$ as transition matrices between various Bethe eigenvectors. We plan to investigate this question further in [KS].
9.3.
In the framework of enumerative geometry, one can define the Frobenius intertwiner as the partition function counting quasimaps from $\mathbb{P}^{1}$ to a Nakajima variety $X$ with two different relative conditions at $0\in\mathbb{P}^{1}$ and $\infty\in\mathbb{P}^{1}$. The corresponding moduli space of quasimaps is represented by Fig.1, we refer to [Oko17] and also [PSZ] where the pictorial notations are explained.
Figure 1. Frobenius intertwiner as a partition function of relative quasimaps
At $0\in\mathbb{P}^{1}$ we have the usual relative condition on quasimaps as defined in [Oko17]. We recall that the relative boundary condition at $0\in\mathbb{P}^{1}$ means that $\mathbb{P}^{1}$ is allowed to have a chain of rational “bubbles” with simple nodal intersections. This chain of bubbles is represented by blue color in Fig. 1.
At $\infty\in\mathbb{P}^{1}$ we also impose a relative condition but require, in addition, the relative quasimaps to be $\mathbb{Z}/p\mathbb{Z}$-invariant. More precisely, recall that the moduli space of relative quasimaps is defined as a stack quotient of quasimaps from a chain of $\mathbb{P}^{1}$’s by the automorphism group (×$(\mathbb{C}^{\times})^{l}$ where $l$ is the length of the chain, see Section 6.2 in [Oko17]. The moduli space of relative quasimaps at $\infty\in\mathbb{P}^{1}$ is constructed as follows: first we consider the space of quasimaps from a chain of rational curves of length $l$ which are invariant with respect to $\mathbb{Z}/p\mathbb{Z}$, which acts in each “bubble” (the component of the chain) via multiplication by $p$ th roots of unity. Then we define the $\mathbb{Z}/p\mathbb{Z}$-invariant quasimaps as the stack quotient of this moduli space by (×$(\mathbb{C}^{\times})^{l}$.
It is known, Theorem 8.1.16 in [Oko17], that the partition function of relative quasimaps from the chain of bubbles at $0\in\mathbb{P}^{1}$ is a fundamental solution matrix $\Psi({\boldsymbol{u}},\hbar,q,z)$ of the quantum difference equation associated with $X$, normalized so that
$\Psi({\boldsymbol{u}},\hbar,q,z)=1+O(z),$
in terminology of [Oko17] $\Psi({\boldsymbol{u}},\hbar,q,z)$ is also referred to as the capping operator.
Similarly, the partition function of $\mathbb{Z}/p\mathbb{Z}$ invariant quasimaps at $\infty\in\mathbb{P}^{1}$ can be shown to be equal to $\Psi({\boldsymbol{u}},\hbar,q^{p},z^{p})^{-1}$. Finally, the contribution of degree 00 - quasimaps which correspond to the “constant” quasimaps from the parameterized component $\mathbb{P}^{1}$ can be computed by $\mathbb{C}^{\times}_{q}$ equivariant localization similarly to how it was done for the shift operators in Section 8.2 of [Oko17], in particular see Lemma 8.2.12. The resulting contribution has the form
$\mathsf{U}(0)=\Gamma_{p,q}\Big{(}(q-\hbar)T^{1/2}X\Big{)},$
where $T^{1/2}X$ denotes the corresponding choice of the polarization bundle of $X$. For instance, in the case $X=T^{*}\mathbb{P}^{1}$ this formula specializes to (3.5). Combining all this together we obtain that the corresponding partition function is
$\mathsf{U}(z)=\Psi({\boldsymbol{u}},\hbar,q,z)\mathsf{U}(0)\Psi({\boldsymbol{u}},\hbar,q^{p},z^{p})^{-1}$
which is exactly the Frobenius intertwiner.
We note that the idea of $\mathbb{Z}/p\mathbb{Z}$ invariant quasimap moduli is similar to the moduli space recently used in [HL24] to define the quantum Steenrod operators in quantum cohomology.
9.4.
Frobenius intertwiners link the theory of $p$ adic differential equations with the theory of exponential sums in number theory [K90]. Namely, interesting sums arise as the traces
(9.1) $\textrm{Tr}(\mathsf{U}(T_{z}))$
where $T_{z}\in\mathbb{Q}_{p}$ denotes the Teichmuller representative of an element $z\in\mathbb{F}_{p}$. These sums may be considered as the arithmetic counterpart of the integral representation for solutions of the corresponding differential equation. See for instance [Dwo74], where the Kloosterman sum is realized as the trace of Frobenius for the Bessel differential equation.
These type of sums for the $q$ case seem never considered. One interesting new feature of the $q$ difference equations for Nakajima varieties is the $3D$-mirror symmetry, which is the invariance of the equations with respect to transformation $z\longleftrightarrow{\boldsymbol{u}}$, where $z$ denote the Kähler parameters (parameters counting the degrees of the quasimaps) and ${\boldsymbol{u}}$ denote the equivariant parameters. It is interesting to investigate what part of this symmetry survives in the arithmetic context of the exponential sums (9.1).
Acknowledgements
I thank K. Kedlaya, J. Hee Lee, A. Oblomkov, D. Vargas-Montoya, M. Vlasenko and especially A. Okounkov for very inspiring discussions. I thank A. Varchenko for introducing me to the subject. This note is the extended version of the talks [Sm23] and [Sm24], I thank the organizers for these opportunities. This work is supported by NSF grants DMS - 2054527 and DMS-2401380.