### v2

# Frobenius intertwiners for $q$-difference equations

Abstract.
We consider a class of $q$-hypergeometric equations describing the quantum difference equation for the cotangent bundles over projective spaces $X=T^{*}\mathbb{P}^{n-1}$ We show that over $\mathbb{Q}_{p}$ these equations are equipped with the Frobenius action $(q,z)\to(q^{p},z^{p})$. We obtain an explicit formula for the constant term of the Frobenius intertwiner in terms of the $p$-adic $q$-gamma function of Koblitz. In the limit $q\to 1$ we arrive at the Frobenius structures for the $p$-adic hypergeometric and Bessel differential equations studied by Dwork. In particular, we find closed formulas for $p$-adic constants appearing in works of Dwork and Sperber in terms of $p$-adic zeta functions.
Department of Mathematics, University of North Carolina at Chapel Hill
Chapel Hill, NC 27599-3250, USA
Key words: Frobenius structures, $q$-hypergeometric functions, $p$-adic $q$-Gamma function, Nakajima varieties, enumerative geometry
2020 Mathematics Subject Classification: 14G33 (11D79, 32G34, 33C05, 33E30)
footnotetext: ${}^{\star}\mskip-0.99998mu$-mail: asmirnov@email.unc.edu
1. Introduction
1.1. $q$-difference equations and enumerative geometry
The quantum K-theory of quiver varieties [Oko17] is governed by the quantum difference equations (QDE) [OS22]. It is expected that over $\mathbb{Q}_{p}$ these equations are equipped with a symmetry which acts on the arguments of the equations by $z\to z^{p}$. In the limit $q\to 1$ these equations reduce to the quantum differential equations of quiver varieties [MO19] while the symmetry $z\to z^{p}$ specializes to the Frobenius structures, which are well known in the theory of $p$-adic differential equations [Dw89, Ked21]. In this sense, the expected symmetry of QDE may be viewed as a $q$-deformation of the Frobenius structures.
One may hope that the $q$-deformed Frobenius structures arising this way may be of significance in arithmetic geometry, see [Sch, BS].
1.2. $q$-hypergeometric case
In this paper consider the $q$-deformed Frobenius structure for simplest examples of quiver varieties given by the cotangent bundles over projective spaces $X=T^{*}\mathbb{P}^{n-1}$. The cotangent bundle $X$ is equipped with a natural action of a torus $T=(\mathbb{C}^{*})^{n+1}$. We denote by $h$ and $\boldsymbol{a}=(a_{1},\dots,a_{n})$ the corresponding equivariant parameters of $T$. The parameter $h$ denotes the character of the one-dimensional representation $\mathbb{C}\omega$ spanned by the symplectic form $\omega\in H^{2}(X,\mathbb{C})$. This parameter plays a special role in the present paper.
In Section 2 we describe the QDE of $X$ which is nothing but the very classical $q$-hypergeometric equation with parameters $h$ and $\boldsymbol{a}$.
1.3. $q$-deformed Frobenius intertwiner
In the context of enumerative geometry it is standard to represent the fundamental solution matrix of QDE as a power series with coefficients in the equivariant $K$-theory:
(1.1) $\Psi(h,\boldsymbol{a},q,z)\in K_{T}(X)^{\otimes 2}[[z]]$
In Section 2 we fix such $\tilde{\Psi}(h,\boldsymbol{a},q,z)$ where $\tilde{}$ refers to a certain specific choice of the normalization, which involves $q$-gamma functions. The key property of this normalization, relevant to our work, is that $\tilde{\Psi}(h,\boldsymbol{a},q,z)$ satisfies a set of regular difference equations in the equivariant parameters $h$ and $\boldsymbol{a}$; see Theorem 3.1.
We consider the power series
(1.2) $\mathsf{U}(h,\boldsymbol{a},q,z)=\tilde{\Psi}(ph,p\boldsymbol{a},q,z)\tilde{\Psi}(h,\boldsymbol{a},q^{p},z^{p})^{-1}\in K_{T}(X)^{\otimes 2}[[z]]$
Fixing a basis in $K_{T}(X)$ we can represent $\mathsf{U}(h,\boldsymbol{a},q,z)$ by an $n\times n$-matrix with coefficients given by power series in $z$. Assume that
$|q-1|_{p}<1,\ \ \textrm{and}\ \ (h,a_{1},\dots,a_{n})\in\mathbb{Z}^{n+1}_{p}$
We show that under these assumptions (Theorem 5.2), $\mathsf{U}(h,\boldsymbol{a},q,z)\in\mathsf{GL}_{n}(\hat{E}_{p})$ where $\hat{E}_{p}$ denotes the field of $p$-adic analytic functions. This means that $\mathsf{U}(h,\boldsymbol{a},q,z)$ is a Frobenius intertwiner between the $q$-hypergeometric systems with parameters $(h,\boldsymbol{a},q^{p},z^{p})$ and $(ph,p\boldsymbol{a},q,z)$.
We observe that the Frobenius intertwiner $\mathsf{U}(h,\boldsymbol{a},q,z)$ has simple specializations at $q$ given by $p$-adic unit roots. If $q$ is a $p$-adic unit root of order $p^{s}$ then
$|q-1|_{p}=p^{-\frac{1}{p^{s-1}(p-1)}}<1$
We show that if $q$ is such a root (Theorem 5.8) then the Frobenius intertwiner $\mathsf{U}(h,\boldsymbol{a},q,z)$ specializes to a rational function of variables $z$ and $q^{ph},\ \ q^{pa_{1}},\dots,q^{pa_{n}}.$ In particular, these rational functions satisfy
$\mathsf{U}(h,\boldsymbol{a},q,z)=\mathsf{U}(h^{(s)},\boldsymbol{a}^{(s)},q,z)$
where $\boldsymbol{a}^{(s)}=(a_{1}^{(s)},\dots,a^{(s)}_{n})$ and $h^{(s)}$ are natural numbers defined by $h^{(s)}=h\pmod{p^{s-1}}$, $a^{(s)}_{i}=a_{i}\pmod{p^{s-1}}$, $i=1,\dots,n$. This property can be considered as a $q$-version of the Dwork’s congruences for the classical hypergeometric functions [Dw69] and the Hasse-Witt matrices [VZ].
1.4. The constant term of the Frobenius intertwiner
In Section 5, we give an explicit formula for the constant term of the Frobenius intertwiner (1.2), i.e., its specialization at $z=0$. First, for a vector bundle $\mathcal{V}$ over $X$ we define a $K$-theory valued characteristic class $\Gamma_{p,q}(\mathcal{V})\in K_{T}(X)$. This class is constructed from the $p$-adic $q$-gamma function $\Gamma_{p,q}$ defined by Koblitz [Ko80]. Second, we show that $\mathsf{U}(h,\boldsymbol{a},q,0)$ is an operator acting in $K_{T}(X)$ as the operator of multiplication by the $K$-theory class
(1.3) $\mathsf{U}(h,\boldsymbol{a},q,0)=\Gamma_{p,q}((q^{p}-q^{ph})T^{1/2}X^{(p)})$
where $T^{1/2}X$ denotes the polarization bundle of $X$ and $T^{1/2}X^{(p)}=\psi^{p}(T^{1/2}X)$ its twist by the $p$-th Adams operation.
For generic values of the equivariant parameters $a_{i}\neq a_{j}$, the equivariant $K$-theory $K_{T}(X)$ has a distinguished basis given by the $K$-theory classes of $T$-fixed points. In this basis the operator $\mathsf{U}(h,\boldsymbol{a},q,0)$ is diagonal, with the eigenvalues
$\mathsf{U}(h,\boldsymbol{a},q,0)_{i,i}=\prod\limits_{j=1}^{n}\,\dfrac{\Gamma_{p,q}(pa_{i}+ph-pa_{j})}{\Gamma_{p,q}(pa_{i}-pa_{j})\Gamma_{p,q}(ph)},\ \ i=1,\dots,n.$
We expect that (1.3) holds for the constant term of the Frobenius intertwiners for more general varieties.
1.5. Limits and specializations
The $q$-deformed Frobenius structure has several important specializations. First, the limit $q\to 1$ corresponds to reduction of the equivariant $K$-theory to the equivariant cohomology. In this limit the QDE specializes to the hypergeometric differential equation - the quantum differential equation of $X=T^{*}\mathbb{P}^{n-1}$. The $p$-adic $q$-gamma function $\Gamma_{p,q}$ specializes to the $p$-adic gamma function $\Gamma_{p}$ of Morita [Mo75]. Combining this together we find that $\left.\mathsf{U}(h,\boldsymbol{a},q,z)\right|_{q=1}$ is the Frobenius structure for the hypergeometric differential equation. Its constant term is the operator acting in the equivariant cohomology $H^{*}_{T}(X)$ as multiplication by the cohomology class $\Gamma_{p}(TX^{(p)})$ where $TX^{(p)}$ denotes the tangent bundle of $X$ twisted by $p$-th Adams operation. For generic values of the equivariant parameters this operator is diagonal in the basis of $T$-fixed points with eigenvalues
(1.4) $\ \ \ \ \ \ \small\mathsf{U}(h,\boldsymbol{a},1,0)_{i,i}=\prod\limits_{w\in\textrm{char}(T_{i}X)}\,\Gamma_{p}(pw)=\prod\limits_{j=1}^{n}\,\Gamma_{p}(pa_{i}+ph-pa_{j})\Gamma_{p}(pa_{j}-pa_{i})\ \ \ i=1,\dots,n.$
where the first product is over the $T$-characters appearing in the tangent space to $X$ at $i$-th $T$-fixed point. In this way we reproduce the results of Dwork [Dw69, Dw89] where the Frobenius intertwiners for the hypergeometric equations were discovered and also give them a geometric interpretation. We also refer to [Ked21] where a formula equivalent to (1.4) was obtained using different tools.
Second, it is well known that in the limit $h\to\infty$ the quantum differential equation of $X=T^{*}\mathbb{P}^{n-1}$ specializes to quantum differential equation of $\mathbb{P}^{n-1}$. In the most degenerate case $a_{1}=a_{2}=\dots=a_{n}$, this equation coincides with the generalized Bessel equation studied over $\mathbb{Q}_{p}$ by Sperber in [Sp80]. For $n=2$ this is the very classical Bessel equation investigated by Dwork [Dw74]. As an application, in the limit $h\to\infty$ we obtain an explicit description for the Frobenius structures of Bessel equations. In particular, we obtain closed formulas for certain $p$-adic constants appearing in [Dw74, Sp80] as values of $p$-adic zeta functions.
1.6. Enumerative definition of Frobenius intertwiner
In quantum cohomology over fields of positive characteristic one can define new classes of operators known as the quantum Steenrod operations [F96, W20]. These operations are defined enumeratively, namely as partition functions counting stable maps from $\mathbb{P}^{1}$ with $p$-marked points to $X$. The stable maps are assumed to be invariant under the cyclic group $\mathbb{Z}/p\mathbb{Z}$, which permutes the marked via rotation of the source $\mathbb{P}^{1}$. It was recently shown in [HL24] that for a large family of varieties, the quantum Steenrod operations coincide with the $p$-curvature of the quantum connection of these varieties.
We expect that this story extends naturally to the level of quantum $K$-theory of quiver varieties which is defined via equivariant count of quasimaps [Oko17]. In particular, the $q$-difference equations have natural analogs of $p$-curvature [KS]. In this context, the quantum Steenrod operations shall be lifted to the $K$-theoretic quantum Adams operations. Moreover, the Frobenius intertwiner $\mathsf{U}(h,\boldsymbol{a},q,z)$ itself can be understood as a partition function counting equivariant quasimaps from $\mathscr{C}=\mathbb{P}^{1}$ to $X$ with two relative boundary conditions at $0,\infty\in\mathscr{C}$, where the relative quasimap moduli space at $\infty$ is assumed to be $\mathbb{Z}/p\mathbb{Z}$-equivariant in an appropriate sense.
When $q$ is specialized at $p$-th unit roots, the partition function $\mathsf{U}(h,\boldsymbol{a},q,z)$ is an operator which conjugates the $p$-curvature of the $q$-difference connection to the Frobenius twist of this connection. In particular, these objects have the same spectrum [KS]. This hints at the arithmetic significance of $q$-deformed Frobenius intertwiners with $q$ specialized at the unit roots.
We briefly touch on these ideas in Section 9.
2. $q$-hypergeometric functions
In this section we work over $\mathbb{C}$. In the next section we switch to the field of $p$-adic numbers.
2.1.
Let $D$ be the $q$-shift operator acting by $Df(z)=f(zq)$. We consider a $q$-difference equation:
(2.1) $P(\boldsymbol{a},\hbar,z)F(z)=0,$
where $P(\boldsymbol{a},\hbar,z)$ denotes the $q$-difference operator:
(2.2) $P(\boldsymbol{a},\hbar,z)=\sum_{m=0}^{n}\,\alpha_{m}(1-z\hbar^{m})D^{m}.$
and $\alpha_{m}$ denotes the coefficient of $x^{m}$ in $\prod\limits_{i=1}^{n}(1-x/u_{i})$.
2.2.
Let us consider the $q$-hypergeometric power series
(2.3) $f_{i}(\hbar,{\boldsymbol{u}},q,z)=\sum\limits_{d=0}^{\infty}\,\Big{(}\prod\limits_{j=1}^{n}\,\dfrac{(u_{i}\hbar/u_{j})_{d}}{(u_{i}q/u_{j})_{d}}\Big{)}\,z^{d},$
where $(x,q)_{d}:=(1-x)(1-xq)\dots(1-xq^{d-1})$. Set $e_{i}(z)=\exp\Big{(}{\frac{\ln(z)\ln(u_{i})}{\ln(q)}}\Big{)},$ so that $e_{i}(zq)=e_{i}(z)u_{i}$. One verifies that the functions
(2.4) ${F}_{i}(\hbar,{\boldsymbol{u}},q,z)=e_{i}(z)f_{i}(\hbar,{\boldsymbol{u}},q,z)$
satisfy (2.1) for all $i=1,\dots,n$. Thus, ${F}_{i}({\boldsymbol{u}},\hbar,q,z)$ give a basis of solutions to (2.1).
2.3.
From (2.1) it is clear that the vectors
(2.9) ${\Psi}_{i}(\hbar,{\boldsymbol{u}},q,z)=\left(\begin{array}[]{l}{F}_{i}(\hbar,{\boldsymbol{u}},q,z)\\ {F}_{i}(\hbar,{\boldsymbol{u}},q,zq)\\ \dots\\ {F}_{i}(\hbar,{\boldsymbol{u}},q,zq^{n-1})\end{array}\right)$
satisfy a first order $q$-difference equation
(2.10) ${\Psi}_{i}(\hbar,{\boldsymbol{u}},q,zq)=\mathbf{M}(\hbar,{\boldsymbol{u}},q,z)\,{\Psi}_{i}(\hbar,{\boldsymbol{u}},q,z)$
for all $i=1,\dots,n$, where $\mathbf{M}(\hbar,{\boldsymbol{u}},q,z)$ is the companion matrix for the characteristic polynomial of (2.2):
$\mathbf{M}(\hbar,{\boldsymbol{u}},q,z)=\left(\begin{array}[]{cccccc}0&1&0&0&\dots&0\\ 0&0&1&0&\dots&0\\ 0&0&0&1&\dots&0\\ ..&..&..&..&..&..\\ \frac{\bar{\alpha}_{n}(z-1)}{1-\hbar^{n}z}&\frac{\bar{\alpha}_{n-1}(z\hbar-1)}{1-\hbar^{n}z}&\frac{\bar{\alpha}_{n-2}(z\hbar^{2}-1)}{\hbar^{n}z-1}&\frac{\bar{\alpha}_{n-3}(z\hbar^{3}-1)}{1-\hbar^{n}z}&\dots&\frac{\bar{\alpha}_{1}(z\hbar^{n-1}-1)}{1-\hbar^{n}z}\\ \end{array}\right)$
where $\bar{\alpha}_{k}$ denotes the coefficients of $x^{k}$ in $\prod_{i=1}^{n}(1-xu_{i})$. The matrix
${\Psi}(\hbar,{\boldsymbol{u}},q,z)=\Big{(}{\Psi}_{1}(\hbar,{\boldsymbol{u}},q,z),\dots,{\Psi}_{n}(\hbar,{\boldsymbol{u}},q,z)\Big{)}.$
with $i$-th column given by ${\Psi}_{i}({\boldsymbol{u}},\hbar,q,z)$ is a fundamental solution matrix of the $q$-hypergeometric system (2.10). Explicitly, it has the following matrix elements
(2.11) ${\Psi}_{i,j}(\hbar,{\boldsymbol{u}},q,z)={F}_{j}(\hbar,{\boldsymbol{u}},q,zq^{i-1}).$
Any matrix with columns given by $\mathbb{C}$ linear combinations of the columns of ${\Psi}({\boldsymbol{u}},\hbar,q,z)$ produces a new fundamental solution matrix. Thus, other (analytic) fundamental solution matrices are of the form
${\Psi}(\hbar,{\boldsymbol{u}},q,z)\Lambda,\ \ \Lambda\in\mathsf{GL}_{n}(\mathsf{\mathbb{C}}).$
Finally, we also note that the fundamental solution matrix has the form
(2.12) ${\Psi}(\hbar,{\boldsymbol{u}},q,z)={\Psi}^{\prime}(\hbar,{\boldsymbol{u}},q,z)\textsf{E}(z).$
where $\textsf{E}(z)$ is the diagonal matrix $\textsf{E}(z)=\operatorname{diag}(e_{1}(z),e_{2}(z),\dots,e_{n}(z))$, and ${\Psi}^{\prime}({\boldsymbol{u}},\hbar,q,z)$ is analytic near $z=0$, i.e., is a power series in $z$.
2.4.
Throughout this paper we will be using both, the multiplicative parameters ${\boldsymbol{u}}=(u_{1},\dots,u_{n})$, $\hbar$ and additive parameters $\boldsymbol{a}=(a_{1},\dots,a_{n})$, $h$ which are related by
(2.13) $u_{1}=q^{a_{1}},\dots,u_{n}=q^{a_{n}},\ \ \hbar=q^{h}.$
In the additive notations we write:
(2.14) $\Psi_{i,j}(h,\boldsymbol{a},q,z)=f_{i}(h,\boldsymbol{a},q,zq^{i-1})z^{a_{j}}q^{a_{j}(i-1)}$
where we denote by $\Psi_{i,j}(h,\boldsymbol{a},q,z)$ and $f_{i}(h,\boldsymbol{a},q,z)$ the functions (2.11) and (2.3) with parameters (2.13) respectively.
3. Vertex function of $X=T^{*}\mathbb{P}^{n-1}$
QDE (2.10) describes the equivariant quantum $K$-theory of the cotangent bundles over projectile spaces [OS22]. We briefly overview the connection in this section.
3.1.
Let (×$(\mathbb{C}^{\times})^{n}$ be the torus acting on $\mathbb{C}^{n}$ by scaling the coordinate lines with characters $u_{1},\dots,u_{n}$. We have an induced action of this torus on the cotangent bundle over projectile space $X=T^{*}\mathbb{P}(\mathbb{C}^{n})$. Let $T=(\mathbb{C}^{\times})^{n}\times\mathbb{C}^{\times}$ be a bigger torus acting on $X$: the first factor (×$(\mathbb{C}^{\times})^{n}$ acts as before and the second $\mathbb{C}^{\times}$ scales the cotangent directions with a character $\hbar^{-1}$. The $T$-equivariant $K$-theory of $X$ is the ring:
(3.1) $K_{T}(X)=\mathbb{Z}[L^{\pm},u_{1}^{\pm},\dots,u^{\pm}_{n},\hbar^{\pm}]/(L-u_{1})\dots(L-u_{n}).$
where $L$ is the class of the line bundle $\mathcal{O}(1)$. Abusing notations we denote by the same symbols $u_{1},\dots,u_{n}$ the trivial equivariant line bundles associated to the characters $u_{i}$. In this language, the additive parameters $a_{i}$, $h$ may be thought of as the first Chern classes of the line bundles $u_{i}$ and $\hbar$. Similarly, let $x$ be the first Chern class of $L$, then in our conventions $L=q^{x}$.
3.2.
Let us define a formal extension of this ring
$\hat{K}_{T}(X)=\mathrm{completion\ of}\ \ K_{T}(X)[z^{x}][[q]],$
so that it contains the following gamma class:
(3.2) $\Phi(h,\boldsymbol{a},x,q,z)=\Big{(}\dfrac{(q^{h},q)_{\infty}}{(q,q)_{\infty}}\Big{)}^{\!\!n}\,\prod\limits_{i=1}^{n}\dfrac{(q^{\,x+1-a_{i}},q)_{\infty}}{(q^{\,x+h-a_{i}},q)_{\infty}}\,z^{\,x}\in\hat{K}_{T}(X),$
where $(u,q)_{\infty}=\prod\limits_{m=0}^{\infty}\,(1-uq^{m})$ is the reciprocal of the $q$ gamma function. The normalized vertex function of $X$ is a generating function counting equivariant quasimaps to quiver varieties, see Section 7.2 of [Oko17] for definitions. For $X$ the vertex function has the following form:
(3.3) $\widetilde{V}(h,\boldsymbol{a},x,q,z)=\sum\limits_{d=0}^{\infty}\,\Phi(h,\boldsymbol{a},x+d,q,z)\,\in\hat{K}_{T}(X)[[z]]$
One verifies that
$P({\boldsymbol{u}},\hbar,z)\,\widetilde{V}(h,\boldsymbol{a},x,q,z)=(L-u_{1})\dots(L-u_{n})\,\Phi(h,\boldsymbol{a},x,q,z)$
where $P({\boldsymbol{u}},\hbar,z)$ is the $q$-difference operator given by (2.2). Note that in (3.1) the vertex function solves the $q$-hypergeometric equation (2.1).
3.3.
The components of the class (3.3) in any chosen basis of $K$-theory (3.1) provide a basis of solutions of the $q$-hypergeometric system (2.1). One convenient basis consists of the $K$-theory classes of the $T$-fixed points $X^{T}$. The components of the class (3.3) in this basis are obtained by the specialization $L=u_{i}$ or, in the additive notations, by the specialization $x=a_{i}$. In this way we obtain the basis of solutions given, up to a normalization multiple, by the hypergeometric series (2.4):
(3.4) $\widetilde{V}(h,\boldsymbol{a},a_{i},q,z)=\Phi(h,\boldsymbol{a},a_{i},q,z){F}_{i}(h,\boldsymbol{a},q,z).$
In this way, the vertex function (3.3) provides a uniform “basis free” description of the solutions to (2.1).
3.4.
It will be convenient to normalize all solutions as in the vertex (3.4). In particular, we introduce the normalized fundamental solution matrix
$\tilde{\Psi}(h,\boldsymbol{a},q,z)=\Big{(}\tilde{\Psi}_{1}(h,\boldsymbol{a},q,z),\dots,\tilde{\Psi}_{n}(h,\boldsymbol{a},q,z)\Big{)},$
with $i$-th column
$\tilde{\Psi}_{i}(h,\boldsymbol{a},q,z)={\Psi}^{\prime}_{i}(h,\boldsymbol{a},q,z)\Phi(h,\boldsymbol{a},a_{i},q,z)$
where ${\Psi}^{\prime}_{i}(h,\boldsymbol{a},q,z)$ is the analytic part of the fundamental solution matrix as defined in (2.12). We can also write the last formula simply as
(3.5) $\tilde{\Psi}(h,\boldsymbol{a},q,z)={\Psi}^{\prime}(h,\boldsymbol{a},q,z)\Phi(h,\boldsymbol{a},x,q,z)$
where $\Phi(\boldsymbol{a},h,x,q,z)$ is now understood as the operator of multiplication by the gamma class (3.2) in the equivariant $K$ theory. In particular, in the basis of the $T$-fixed points this operator is represented by the diagonal matrix with eigenvalues given by $\Phi(h,\boldsymbol{a},a_{i},q,z)$, $i=1,\dots,n$.
3.5.
The normalized fundamental solution matrix $\tilde{\Psi}(\boldsymbol{a},h,q,z)$ has an enumerative meaning: it coincides with the capping operator for $X$, which is the partition function of the relative quasimaps to $X$, see Section 7.4 in [Oko17] for the definitions. In particular, normalized as in (3.5) the fundamental solution matrix satisfies the shift equations which are the difference equations in the equivariant parameters $\boldsymbol{a}$ and $h$. These equations exist in general for an arbitrary Nakajima variety. Applied to $X=T^{*}\mathbb{P}^{n-1}$ they can be formulated as follows:
Theorem 3.1 (Theorem 8.2.20, [Oko17]).
The normalized fundamental solution matrix (3.5) satisfies the following system of difference equations
$\tilde{\Psi}(h,a_{1},\dots,a_{i}+1,\dots a_{n},q,z)=A_{i}(\boldsymbol{a},h,q,z)\tilde{\Psi}(h,a_{1},\dots,a_{i},\dots a_{n},q,z),$
$\tilde{\Psi}(h+1,a_{1},\dots,a_{i},\dots a_{n},q,z)=H(\boldsymbol{a},h,q,z)\tilde{\Psi}(h,a_{1},\dots,a_{i},\dots a_{n},q,z).$
where $A_{i}(\boldsymbol{a},h,q,z)$ and $H(\boldsymbol{a},h,q,z)$ denote matrices whose entries are rational functions in the equivariant parameters ${\boldsymbol{u}}$, $\hbar$, $q$ and $z$: $A_{i}(\boldsymbol{a},h,q,z),H(\boldsymbol{a},h,q,z)\in\mathsf{GL}_{n}(\mathbb{Q}({\boldsymbol{u}},\hbar,z,q)).$
These difference equations are useful for the analysis of QDEs, particularly when the parameters $h$, $\boldsymbol{a}$ are specialized to $\mathbb{Z}_{p}$-values, to which we now switch out attention.
4. $p$ adic completions and norms
In this section, we provide a brief overview of $p$-adic analysis, $p$-adic analytic functions, and give several illustrative examples of simple Frobenius structures. Our running example in this section is the quantum differential and quantum difference equation associated with a “point” $X=T^{*}\mathbb{P}^{0}$. The Frobenius structures arising in this case can be described using only elementary combinatorics.
4.1.
We denote by $|\cdot|_{p}$ the $p$-adic norm on $\mathbb{Q}$ normalized so that $|p^{s}|=1/p^{s}$. The field of $p$-adic numbers $\mathbb{Q}_{p}$ is the completion of $\mathbb{Q}$ with respect to this norm. $\mathbb{Z}_{p}\subset\mathbb{Q}_{p}$ denotes the set of integers of $\mathbb{Q}_{p}$ which is the subset of elements with norm $|x|_{p}\leqslant 1$. We also denote by $\Omega$ the completion of the algebraic closure of $\mathbb{Q}_{p}$.
Let $f(z)$ be a rational function:
$f(z)=\dfrac{\sum_{i}a_{i}z^{i}}{\sum_{j}b_{j}z^{j}}\in\mathbb{Q}(z)$
The $p$-adic Gauss norm on $\mathbb{Q}(z)$ is defined by
$|f(z)|_{\rm Gauss}:=\dfrac{\textrm{max}_{i}\{|a_{i}|_{p}\}}{\textrm{max}_{j}\{|b_{i}|_{p}\}}$
We denote by $E_{p}$ the field of $p$-adic analytic functions which is the completion of $\mathbb{Q}(z)$ with respect to the Gauss norm. By definition, an element $f(z)\in E_{p}$ is a sum
$f(z)=\sum\limits_{n=0}^{\infty}\,f_{i}(z),\ \ f_{i}(z)\in\mathbb{Q}(z)$
converging in the Gauss norm. Note that $|f(z)|_{\rm Gauss}=\max\{|f_{i}(z)|_{\rm Gauss}:i=0,1,\dots\}.$
4.2.
One of the main questions we encounter in this paper is the following: let $f(z)\in\mathbb{Q}[[z]]$ be a power series. When is $f(z)$ a Taylor expansion of an element $g(z)\in E_{p}$?
Since the norm of analytic elements is bounded, there exist $n$ such that $|p^{n}g(z)|_{\rm Gauss}\leqslant 1$. Thus, if $f(z)$ is the series expansion of $g(z)\in E_{p}$ then we can “clear denominators” - the coefficients of power series $p^{n}f(z)$ do not have powers of $p$ appearing in the denominators. In particular, the coefficients of the power series $p^{n}f(z)$ have well defined reductions $\pmod{p^{s}}$ for $s=1,2,\dots$. These reductions can be used to determine whether $f(z)$ is the expansion of an analytic function using the following elementary lemma.
Lemma 4.1.
A power series $f(z)\in\mathbb{Q}[[z]]$ is a Taylor expansion of an analytic element $g(z)\in E_{p}$ if and only if there exist $n\in\mathbb{N}$ such that for any $s\in\mathbb{N}$
(4.1) $p^{n}f(z)\equiv g_{s}(z)\pmod{p^{s}}$
where $g_{s}(z)$ is a power series expansion of some rational function from $\mathbb{Q}(z)$.
Proof.
First assume $f(z)\in\mathbb{Q}[[z]]$ is such that (4.1) holds for some $n$. Let us define $f_{i}(z)$ by
$f_{i}(z)=g_{i+1}(z)-g_{i}(z)$
By construction $|f_{i}(z)|_{p}<1/p^{i}$, thus the sum $\sum_{i=0}^{\infty}f_{i}(z)$ converges in the Gauss norm and we obtain:
$f(z)=\dfrac{1}{p^{n}}\sum_{i=0}^{\infty}f_{i}(z)\in E_{p}$
For the opposite direction, assume $f(z)$ is a power series in $z$ which is a Taylor expansion of some ${g}(z)\in E_{p}$. We have
(4.2) ${g}(z)=\sum\limits_{i=0}^{\infty}\,g_{i}(z)$
for some $g_{i}(z)\in\mathbb{Q}(z)$. Assume that $|{g}(z)|_{\rm Gauss}=p^{n}$ for some $n\in\mathbb{Z}$. Since (4.2) converges in the Gauss norm it follows that for any $s\in\mathbb{N}$ there is $N_{s}\in\mathbb{N}$ such that $p^{n}g_{i}(z)\equiv 0\pmod{p^{s}}$ for all $i>N_{s}$. Therefore,
$p^{n}f(z)\equiv\sum\limits_{n=0}^{N_{s}}\,g_{i}(z)\pmod{p^{s}}$
In Section 8 we will demonstrate how this Lemma can be used to discover non-trivial Frobenius structures numerically.
Example: Let $h\in\mathbb{Q}$ be a rational number such that $h\in\mathbb{Z}_{p}$. Let us consider a power series
(4.3) $f(h,z)=\sum\limits_{n=0}^{\infty}\,(-1)^{n}\binom{h}{n}z^{n}$
which is a Taylor expansion of $(1-z)^{h}$. It is known that this series is not an expansion of an element from $E_{p}$ for generic $h$. However, the power series
(4.4) $\mathsf{U}(h,z)=\dfrac{f(ph,z)}{f(h,z^{p})}$
is in $E_{p}$. Indeed
$\mathsf{U}(h,z)=\dfrac{(1-z)^{ph}}{(1-z^{p})^{h}}=\Big{(}1+\dfrac{(1-z)^{p}-(1-z^{p})}{1-z^{p}}\Big{)}^{h}$
Note that $(1-z)^{p}-(1-z^{p})=p\alpha(z)$ for some $\alpha(z)\in\mathbb{Z}[z]$ and therefore
$\mathsf{U}(h,z)=\sum\limits_{n=0}^{\infty}\,\binom{h}{n}p^{n}\dfrac{\alpha(z)^{n}}{(1-z^{p})^{n}}$
For $h\in\mathbb{Z}_{p}$ we also have $\binom{h}{n}\in\mathbb{Z}_{p}$ for any $n$. From which we see that:
(4.5) $\mathsf{U}(h,z)\equiv\sum\limits_{n=0}^{s-1}\,\binom{h}{n}p^{n}\dfrac{\alpha(z)^{n}}{(1-z^{p})^{n}}\pmod{p^{s}}$
Therefore $\mathsf{U}(h,z)\in E_{p}$ by the previous Lemma.
4.3.
Let $V(\alpha,z)$ denote the finite dimensional $\mathbb{Q}_{p}$-space of solutions to some $p$-adic differential equation in $z$ with parameters $\alpha$. As we explain in the next section, a Frobenius intertwiner gives a map:
$V(\alpha,z)\stackrel{{\scriptstyle\mathsf{U}(z)}}{{\longrightarrow}}V(\alpha^{\prime},z^{p})$
As a map between two different vector spaces $\mathsf{U}(z)$ carries no interesting information. However, when $\alpha^{\prime}=\alpha$ at the points $z^{p}=z$ it becomes an automorphism of $V(\alpha,z)$ and the characteristic polynomial of $\mathsf{U}(z)$ at such points is an interesting invariants of the QDEs. The invariants appearing in this way find many applications in the theory of Gauss and exponential sums [Dw74, Sp80, K90].
The points $z\in\mathbb{Q}_{p}$ satisfying $z^{p}=z$ are the Teichmüller elements: recall that for any $a\in\mathbb{F}_{p}$ there is unique $[a]\in\mathbb{Q}_{p}$ such that $[a]^{p}=[a]$ and $[a]\equiv a\pmod{p}$. Clearly, the map $\mathbb{F}_{p}^{\times}\to\mathbb{Q}^{\times}_{p}$, sending $a$ to $[a]$, is a multiplicative character. The element $[a]\in\mathbb{Q}_{p}$ is called the Teichmüller representative or the Teichmüller lift of $a$.
By definition, if $a\neq 0$ then $|[a]|_{p}=1$. For this reason, a power series $f(z)\in\mathbb{Q}[[z]]$ usually can not be evaluated at such points, since the convergence is not guaranteed. A notable property of the analytic elements $g(z)\in E_{p}$ is that they often can be evaluated at points $z\in\mathbb{Z}_{p}$ with $|z|_{p}=1$: if $f(z)\in E_{p}$ then up to terms with norm $1/p^{s}$ it can be approximated by a rational function. The values of these rational functions at $[a]$ then approximate $f([a])$ with an error of norm $1/p^{s}$.
For these reasons, Frobenius intertwiners are naturally required to have coefficients in $E_{p}$. This motivates the definition we give in the next section.
Example: Continuing our running example, we note that the power series (4.3) spans $1$-dimensional space of solutions of the simplest hypergeometric differential equation
(4.6) $(1-z)z\dfrac{df(z)}{dz}+zhf(z)=0$
In the hypergeometric notations $f(z)={}_{0}F_{1}(-h;;z)$. By its definition, multiplication by power series (4.4) maps solutions with parameters $(h,z^{p})$ to solutions with parameters $(ph,z)$. As we have already seen, $U(h,z)$ is in $E_{p}$. Thus, $U(h,z)$ is a Frobenius intertwiner between these solutions.
Assume that $h$ is such that $(p-1)h\in\mathbb{Z}$. Then $(1-z)^{h(1-p)}$ is a rational function and therefore the composition
(4.7) $\mathsf{U}^{\prime}(z)=(1-z)^{h(1-p)}\mathsf{U}(h,z)=\dfrac{(1-z)^{h}}{(1-z^{p})^{h}}$
is also in $E_{p}$. Thus $\mathsf{U}^{\prime}(z)$ is a Frobenius intertwiner between solutions at $(h,z^{p})$ and $(h,z)$. Using the approximations by rational functions (4.5) one can prove that the value of this intertwiner at the Teichmüller elements equals:
(4.8) $\mathsf{U}^{\prime}([a])=[a-1]^{h(1-p)}$
Note that a naive substitution of $z=[a]$ to (4.7) would give an incorrect answer $\mathsf{U}^{\prime}([a])=1$ (since $[a]^{p}=[a]$). The $p$-adic unit root (4.8) is the “interesting invariant” of the differential equation (4.6) which we mentioned above.
4.4.
Let show that $\mathsf{U}(h,z)$ given by (4.4) is in $E_{p}$ in a different way using the $q$-deformation. First, let us fix $q\in\Omega$ of the form $q=1+t$ with $|t|_{p}<1$. In this case, for $h\in\mathbb{Z}_{p}$ we have well defined elements $q^{\pm h}\in\Omega$. The $q$-deformation of the hypergeometric $f(z)={}_{0}F_{1}(-h;;z)$ is given by the power series
(4.9) $f(h,q,z)=\sum\limits_{n=0}^{\infty}\,\dfrac{(q^{-h},q)_{n}}{(q,q)_{n}}z^{n},\ \ \ (x,q)_{n}=(1-x)(1-xq)\dots(1-xq^{n-1}),$
which solves the $q$-analog of (4.6):
(4.10) $f(h,q,zq)=\dfrac{(1-z)}{(1-zq^{-h})}f(h,q,z)$
At $h\in\mathbb{N}$ this power series truncates to a Laurent polynomial
(4.11) $f(h,q,z)=(1-zq^{-1})(1-zq^{-2})\dots(1-zq^{-h}).$
Assume that $\zeta\neq 1$ is a unit root $\zeta^{p^{s}}=1$. Assume further that the order of the root $\zeta$ is $p^{m}$ for some $m\leqslant s$. Let $h\in\mathbb{N}$ and write $h=h^{(s)}+p^{s-1}h^{\prime}$ for some natural numbers $h^{(s)}$ and $h^{\prime}$. From (4.11) we find
$f(ph,\zeta,z)=f(ph^{(s)},\zeta,z)(1-z^{p^{m}})^{h^{\prime}p^{s-m}}$
Since $\zeta^{p}$ is also a unit root or order $p^{m-1}$ we obtain similarly
$f(h,\zeta^{p},z^{p})=f(h^{(s)},\zeta^{p},z^{p})(1-z^{p^{m}})^{h^{\prime}p^{s-m}}$
where in the last two equalities we used that $\prod_{i=0}^{m-1}(1-z\zeta^{i})=(1-z^{m})$ if $\zeta$ has order $m$.
Let us consider the power series
$\mathsf{U}(h,q,z)=\dfrac{f(ph,q,z)}{f(h,q^{p},z^{p})}$
From the above computations, we see that for $h\in\mathbb{N}$ we have
(4.12) $\mathsf{U}(h,\zeta,z)=\mathsf{U}(h^{(s)},\zeta,z)$
where $h^{(s)}$ is a natural number defined by $h^{(s)}=h\pmod{p^{s-1}}$.
The equation (4.12) is crucial. Note that (4.12) extends to $p$-adic integers $h\in\mathbb{Z}_{p}$. Clearly, unit roots $q^{p^{s}}=1$ such that $q\neq 1$ are the roots of the polynomial:
$[p^{s}]_{q}=\dfrac{1-q^{p^{s}}}{1-q}$
Thus, the relation (4.12) implies that for any $h\in\mathbb{Z}_{p}$ we have
$\mathsf{U}(h,q,z)=\mathsf{U}(h^{(s)},q,z)+[p^{s}]_{p}\,\Big{(}\ \ \dots\ \ \Big{)}$
where $\mathsf{U}(h^{(s)},q,z)$ is a rational function (since $h^{(s)}$ is integral) and $\dots$ stands for a power series in $z$ with coefficients regular at $p^{s}$-th roots. Note that for $q=1+t$ with $|t|_{p}<1$ we have $[p^{l}]_{p}=0\pmod{p^{s}}$ for sufficiently large $l$ thus
(4.13) $\mathsf{U}(h,q,z)\pmod{p^{s}}.$
is a rational function of $z$.
In the limit $q\to 1$ $\mathsf{U}(h,q,z)$ converges to (4.4) and therefore $\mathsf{U}(h,z)\pmod{p^{s}}.$ is also rational and by Lemma 4.1 we conclude that $\mathsf{U}(h,z)\in E_{p}$.
The above computation (4.13) also shows what $\mathsf{U}(h,q,z)\in\hat{E}_{p}$ where $\hat{E}_{p}$ is the completion of $\Omega(z)$ with respect to Gauss norm (we need to replace $\mathbb{Q}$ with $\Omega$ to accommodate for $q\in\Omega$). Thus, $\mathsf{U}(h,q,z)$ is a Frobenius intertwiner between $f(h,q^{p},z^{p})$ and $f(ph,q,z)$.
Finally, let us note that the hypergeometric differential equation (4.6) is the quantum differential equation associated with a zero-dimensional quiver variety $X=T^{*}\mathbb{P}^{0}$ (enumerative geometry of such varieties may be non-trivial, see [DS]). The $q$-difference equation (4.10) is the $K$-theoretic quantum difference equation for this $X$. The power series (4.3) and (4.9) represent the “fundamental solution matrices” of these equations. The running example of this section is thus, the most basic $n=1$ case of the hypergeometric functions (2.3) we consider in this paper.
5. Frobenius intertwiner for $q$-hypergeometric equation
5.1.
Let us consider the $q$-difference hypergeometric equation (2.10). Unless otherwise specified, we assume throughout this section that $h\in\mathbb{Z}_{p}$, $\boldsymbol{a}\in\mathbb{Z}^{n}_{p}$. We assume that $q\in\Omega$ is of the form $q=1+t$ with $|t|_{p}<1$. In this case
$\hbar=q^{h},\ \ u_{1}=q^{a_{1}},\dots,u_{n}=q^{a_{n}}$
are well defined in $\Omega$. Let $\hat{E}_{p}$ be the completion of $\Omega(z)$ with respect to the Gauss norm. The following is motivated by discussion of Section 4.
Definition 5.1.
A Frobenius intertwiner from a $q$-hypergeometric system with parameters $(h,\boldsymbol{a},q^{p},z^{p})$ to a $q$-hypergeometric system with parameters $(h^{\prime},\boldsymbol{a}^{\prime},q,z)$ is an element $\mathsf{U}(z)\in\mathsf{GL}_{n}(\hat{E}_{p})$ such that $\mathbf{M}(\hbar,\boldsymbol{a},q^{p},z^{p})=\mathsf{U}(zq)^{-1}\,\mathbf{M}(\hbar^{\prime},\boldsymbol{a}^{\prime},q,z)\,\mathsf{U}(z)$.
If ${\Psi}(h,\boldsymbol{a},q^{p},z^{p})$ is a fundamental solution matrix for the $q$-hypergeometric system with parameters $(h,\boldsymbol{a},q^{p},z^{p})$ and $\mathsf{U}(h,\boldsymbol{a},q,z)$ is a Frobenius intertwiner as in the Definition above then $\mathsf{U}(h,\boldsymbol{a},q,z){\Psi}(h,\boldsymbol{a},q^{p},z^{p})$ is a fundamental solution matrix for the $q$-hypergeometric system with parameters $(h^{\prime},\boldsymbol{a}^{\prime},q,z)$. Thus, it is of the form ${\Psi}(h^{\prime},\boldsymbol{a}^{\prime},q,z)\Lambda$ where $\Lambda\in\mathsf{GL}_{n}(\Omega)$ is some constant matrix. We obtain a relation
(5.1) ${\Psi}(h^{\prime},\boldsymbol{a}^{\prime},q,z)\Lambda=\mathsf{U}(h,\boldsymbol{a},q,z){\Psi}(h,\boldsymbol{a},q^{p},z^{p}).$
$\mathsf{U}(h,\boldsymbol{a},q,0)$ is called the constant term of the Frobenius intertwiner. If the fundamental solution matrix is normalized by ${\Psi}(h,\boldsymbol{a},q,0)={\rm Id}$ (the identity matrix), then $\mathsf{U}(0)=\Lambda$.
5.2.
Let us consider the power series:
(5.2) $\mathsf{U}(h,\boldsymbol{a},q,z)=\tilde{\Psi}(ph,p\boldsymbol{a},q,z)\tilde{\Psi}(h,\boldsymbol{a},q^{p},z^{p})^{-1}\in K_{T}(X)^{\otimes 2}[[z]]$
where $\tilde{\Psi}(h,\boldsymbol{a},q,z)$ denotes the fundamental solution matrix normalized as in (3.5). Using the natural non-degenerate paring on $K_{T}(X)$ we can also view it as an operator
$\mathsf{U}(h,\boldsymbol{a},q,z):K_{T}(X)[[z]]\longrightarrow K_{T}(X)[[z]]$
In particular, the constant term is a linear map in $K$-theory
(5.3) $\mathsf{U}(h,\boldsymbol{a},q,0):K_{T}(X)\longrightarrow K_{T}(X)$
The goal of this section is to prove the following result (see Section 3.1 for notations and description of the ring $K_{T}(X)$):
Theorem 5.2.
The power series $\mathsf{U}(h,\boldsymbol{a},q,z)$ defined by (5.2) is a Frobenius intertwiner between the $q$-hypergeometric systems with parameters $(h,\boldsymbol{a},q^{p},z^{p})$ and $(ph,p\boldsymbol{a},q,z)$, where $p\boldsymbol{a}=(pa_{1},\dots,pa_{n})$. The constant term of the intertwiner (5.3) acts in $K_{T}(X)$ as the operator of multiplication by the K-theory class
(5.4) $\mathsf{U}(h,\boldsymbol{a},q,0)=\prod\limits_{j=1}^{n}\,\dfrac{\Gamma_{p,q}(px+ph-pa_{j})}{\Gamma_{p,q}(px-pa_{j})\Gamma_{p,q}(ph)}.$
Definition of $p$-adic $q$-Gamma function $\Gamma_{p,q}$ is reminded in Section A.3.
5.3.
Let us first analyze the constant term of this operator.
Proposition 5.3.
The constant term of the power series (5.2) acts in $K_{T}(X)$ as multiplication by the following K-theory class:
(5.5) $\mathsf{U}(h,\boldsymbol{a},q,0)=\prod\limits_{j=1}^{n}\,\dfrac{\Gamma_{p,q}(px+ph-pa_{j})}{\Gamma_{p,q}(px-pa_{j})\Gamma_{p,q}(ph)}$
where $\Gamma_{p,q}$ denotes the $p$-adic $q$-gamma function .
Proof.
From the normalization (3.5) we find:
$\mathsf{U}(h,\boldsymbol{a},q,0)=\dfrac{\Phi(ph,p\boldsymbol{a},px,q,z)}{\Phi(h,\boldsymbol{a},x,q^{p},z^{p})}$
From definition (3.2) we see that this ratio is exactly the product of factors as in (A.7). The Proposition follows by Lemma A.4. ∎
If the set of the $T$-fixed points $X^{T}$ is finite, then multiplication by any equivariant $K$-theory class is diagonal in the basis of $K_{T}(X)$ given by the classes of $T$-fixed points. In our case $X=T^{*}\mathbb{P}^{n-1}$ the set $X^{T}$ is finite when all $a_{1},\dots,a_{n}$ are pairwise distinct. The corresponding eigenvalues of multiplication by $\mathsf{U}(h,\boldsymbol{a},q,0)$ are obtained by specialization $x=a_{i}$. We thus conclude with the following result.
Corollary 5.4.
If $a_{1},\dots,a_{n}$ are pairwise distinct then the constant term of the Frobenius intertwiner $\mathsf{U}(h,\boldsymbol{a},q,0)$ is diagonal in the basis of torus fixed points of $K_{T}(X)$. The corresponding eigenvalues are equal:
(5.6) $\mathsf{U}(h,\boldsymbol{a},q,0)_{i,i}=\prod\limits_{j=1}^{n}\,\dfrac{\Gamma_{p,q}(p(a_{i}-a_{j}+h))}{\Gamma_{p,q}(p(a_{i}-a_{j}))\Gamma_{p,q}(ph)},\ \ \ i=1,\dots,n.$
This result has the following $K$-theoretic formulation. Let $\mathcal{V}$ be a rank $r$ vector bundle over $X$. Using the splitting principle, we may assume that in $K$-theory $\mathcal{V}$ is a sum of $r$ line bundles with Chern roots $x_{1},\dots,x_{r}$. We can thus define a $K$-theory valued characteristic class by
(5.7) $\Gamma_{p,q}(\mathcal{V})=\prod_{i=1}^{r}\Gamma_{p,q}(x_{i})^{-1}$
Recall that every Nakajima variety is equipped with a well-defined polarization bundle $T^{1/2}X$, i.e., $T^{1/2}X$ is a $K$-theory class representing a “half of the tangent bundle” [MO19]:
$TX=T^{1/2}X+\hbar^{-1}\overline{T^{1/2}X}$
where $\hbar$ is the $T$-character of the symplectic form. In our case $X=T^{*}\mathbb{P}^{n-1}$, the polarization can be chosen in the form $T^{1/2}X=W^{*}\otimes L-1$. Where $W$ is a trivial rank $r$ vector bundle with Chern roots given by the equivariant parameters $a_{1},\dots,a_{n}$, and $L$ is the tautological line bundle with Chern root $x$. We note that the arguments of the gamma functions in (5.4) is over the $p$-multiples of the Chern roots of the virtual bundle $(q-\hbar)T^{1/2}X$, in other words is over the Chern roots of $\psi^{p}((q-\hbar)T^{1/2}X)$ where $\psi^{p}$ is the $p$-th Adams operation in the equivariant $K$-theory (which acts on the Chern roots by $x_{i}\to px_{i}$). Combining all this together we obtain:
Proposition 5.5.
The constant term of (5.3) acts as multiplication by the K-theory class
(5.8) $\mathsf{U}(h,\boldsymbol{a},q,0)=\Gamma_{p,q}((q^{p}-\hbar^{p})T^{1/2}X^{(p)})$
where $T^{1/2}X^{(p)}$ is the polarization bundle of $X$ twisted by the $p$-th Adams operation.
We expect that in this form the Proposition holds for every Nakajima variety $X$.
5.4.
Let us consider the vertex function (3.3) normalized by a factor $\Phi(\boldsymbol{a},h,x,q,z)$ given by (3.2). If $h\in\mathbb{N}$, using the notations as in (A.3), (A.4) we can write the vertex function in the form
$\widetilde{V}(h,\boldsymbol{a},x,q,z)=z^{x}\,\sum\limits_{d=0}^{\infty}\,\Big{(}\prod\limits_{i=1}^{n}\dfrac{[q^{\,x+d-a_{i}},q]_{h}}{[1,q]_{h}}\Big{)}z^{d}.$
Proposition 5.6.
Let $\zeta\neq 1$ be a unit root $\zeta^{p^{s}}=1$ and let $h^{\prime}\in\mathbb{N}$ then we have the following identity for the vertex functions
$\left.\tilde{V}(h^{\prime}p^{s},p\boldsymbol{a},px,q,z)\right|_{q=\zeta}=\left.\tilde{V}(h^{\prime}p^{s-1},\boldsymbol{a},x,q^{p},z^{p})\right|_{q=\zeta}$
Proof.
We have:
$\left.\tilde{V}(h^{\prime}p^{s},p\boldsymbol{a},px,q,z)\right|_{q=\zeta}=z^{px}\,\sum\limits_{d=0}^{\infty}\,\left(\prod\limits_{i=1}^{n}\left.\dfrac{[q^{px-pa_{i}+d},q]_{h^{\prime}p^{s}}}{[1,q]_{h^{\prime}p^{s}}}\right)\right|_{q=\zeta}z^{d}$
By Lemma A.1 the coefficients of this sum vanish unless $p\mid d$, therefore we write
$\left.\tilde{V}(h^{\prime}p^{s},p\boldsymbol{a},px,q,z)\right|_{q=\zeta}=z^{px}\,\sum\limits_{d=0}^{\infty}\,\left(\prod\limits_{i=1}^{n}\left.\dfrac{[q^{px-pa_{i}+pd},q]_{h^{\prime}p^{s}}}{[1,q]_{h^{\prime}p^{s}}}\right)\right|_{q=\zeta}z^{pd}$
again, by Lemma A.1 we obtain
$z^{px}\,\sum\limits_{d=0}^{\infty}\,\left.\left(\prod\limits_{i=1}^{n}\dfrac{[q^{px-pa_{i}+pd},q]_{h^{\prime}p^{s}}}{[1,q]_{h^{\prime}p^{s}}}\right)\right|_{q=\zeta}z^{pd}=z^{px}\,\sum\limits_{d=0}^{\infty}\,\left.\left(\prod\limits_{i=1}^{n}\dfrac{[q^{px-pa_{i}+pd},q^{p}]_{h^{\prime}p^{s-1}}}{[1,q^{p}]_{h^{\prime}p^{s-1}}}\right)\right|_{q=\zeta}z^{pd}$
and the right side is $\left.\tilde{V}(h^{\prime}p^{s-1},\boldsymbol{a},x,q^{p},z^{p})\right|_{q=\zeta}$. ∎
Since the $K$-theory components of the vertex functions provide a basis in the space of solutions for the quantum difference equation, the above Proposition extends to the fundamental solution matrix which is normalized as the vertex function (3.3), i.e., the fundamental solution matrix (3.5). Thus, we have:
Theorem 5.7.
The ratio of fundamental solution matrices (5.2)
$\widetilde{\Psi}(h^{\prime}p^{s},p\boldsymbol{a},q,z)\widetilde{\Psi}(h^{\prime}p^{s-1},\boldsymbol{a},q^{p},z^{p})^{-1}$
is well defined at $q$ given by unit roots $\zeta\neq 1$, $\zeta^{p^{s}}=1$ and we have
(5.9) $\left.[\widetilde{\Psi}(h^{\prime}p^{s},p\boldsymbol{a},q,z)\widetilde{\Psi}(h^{\prime}p^{s-1},\boldsymbol{a},q^{p},z^{p})^{-1}\right]_{q=\zeta}=1.$
Proof.
For $h^{\prime}\in\mathbb{N}$ the Theorem follows from the Previous proposition. Since the right side of (5.9) is independent of $h^{\prime}$ and $\mathbb{N}$ is dense in $\mathbb{Z}_{p}$, this identity extends to $h^{\prime}\in\mathbb{Z}_{p}$. ∎
5.5.
We note that the proof of Theorem 5.7 uses only the fact that the $K$-theory components of the vertex function generate the whole $q$-holonomic module, see for instance Theorem 4 in [AO21]. For this reason, this argument applies to any choice of the equivariant parameters including the most degenerate case $a_{i}=0$, which corresponds to the limit to the non-equivariant $K$-theory ring $K(X)$.
In the non-degenerate case $a_{i}\neq a_{j}$, $K_{T}(X)$ has a distinguished basis given by classes of the torus fixed points. In this basis the components of the fundamental solution matrix $\widetilde{\Psi}(p\boldsymbol{a},h^{\prime}p^{s},q,z)$ are given explicitly by $q$-hypergeometric series. We can use this presentation to prove Theorem 5.7 using only elementary combinatorics. For future references, we outline such a proof in this section.
Assume $a_{i}\neq a_{j}$. First, from normalization (3.5) we have
$\widetilde{\Psi}(h^{\prime}p^{s},p\boldsymbol{a},q,z)\widetilde{\Psi}(h^{\prime}p^{s-1},\boldsymbol{a},q^{p},z^{p})^{-1}={\Psi}(h^{\prime}p^{s},p\boldsymbol{a},q,z)\,\mathsf{U}(h^{\prime}p^{s-1},\boldsymbol{a},q,0)\,{\Psi}(h^{\prime}p^{s-1},\boldsymbol{a},q^{p},z^{p})^{-1}$
In the basis of $T$-fixed points in $K_{T}(X)$, ${\Psi}(\boldsymbol{a},h,q,z)$ has matrix elements (2.14). The middle term is the multiplication by $K$-theory class (5.5). In the basis of $T$-fixed this operator is diagonal with the eigenvalues given by:
(5.10) $\mathsf{U}(h^{\prime}p^{s-1},\boldsymbol{a},q,0)_{i,i}=\prod\limits_{j=1}^{n}\,\dfrac{\Gamma_{p,q}(pa_{i}+h^{\prime}p^{s}-pa_{j})}{\Gamma_{p,q}(pa_{i}-pa_{j})\Gamma_{p,q}(h^{\prime}p^{s})}$
At $q$ given by $p^{s}$-th unit roots $\mathsf{U}(h^{\prime}p^{s-1},\boldsymbol{a},q,0)=1$ from Proposition A.5.
Next, from (2.14), the matrix elements of the fundamental solution matrix have the form
(5.11) ${\Psi}_{i,j}(h^{\prime}p^{s},p\boldsymbol{a},q,z)=f_{i}(h^{\prime}p^{s},p\boldsymbol{a},q,zq^{i-1})z^{pa_{j}}q^{pa_{j}(i-1)}$
and similarly
(5.12) ${\Psi}_{i,j}(h^{\prime}p^{s-1},\boldsymbol{a},q^{p},z^{p})=f_{i}(h^{\prime}p^{s-1},\boldsymbol{a},q^{p},z^{p}q^{p(i-1)})z^{pa_{j}}q^{pa_{j}(i-1)}$
If $\zeta\neq 1$ and $\zeta^{p^{s}}=1$ we can assume that $\zeta$ is a primitive $p^{m}$-th unit root for some $1\leqslant m\leqslant s$, then, by part b) of Lemma A.2 we obtain
$\left.f_{i}(h^{\prime}p^{s},p\boldsymbol{a},q,zq^{i-1})\right|_{q=\zeta}=\left.f_{i}(h^{\prime}p^{s-m}p^{m},p\boldsymbol{a},q,zq^{i-1})\right|_{q=\zeta}=(1-z^{p^{m}})^{p^{s-m}h^{\prime}}$
Since $\zeta^{p}$ is also a primitive $p^{m-1}$-th unit root we obtain similarly:
$\left.f_{i}(h^{\prime}p^{s-1},\boldsymbol{a},q^{p},z^{p}q^{p(i-1)})\right|_{q=\zeta}=\left.f_{i}(h^{\prime}p^{s-m}p^{m-1},\boldsymbol{a},q^{p},z^{p}q^{p(i-1)})\right|_{q=\zeta}=(1-z^{p^{m}})^{p^{s-m}h^{\prime}}$
Therefore the matrices (5.11) and (5.12) are equal:
$\left.{\Psi}_{i,j}(h^{\prime}p^{s},p\boldsymbol{a},q,z)\right|_{q=\zeta}=\left.{\Psi}_{i,j}(h^{\prime}p^{s-1},\boldsymbol{a},q^{p},z^{p})\right|_{q=\zeta}=(1-z^{p^{m}})^{p^{s-m}h^{\prime}}q^{pa_{j}(i-1)}$
Note also that for generic values $a_{i}\neq a_{j}$ these matrices are invertible and therefore
$\left.[\widetilde{\Psi}(h^{\prime}p^{s},p\boldsymbol{a},q,z)\widetilde{\Psi}(h^{\prime}p^{s-1},\boldsymbol{a},q^{p},z^{p})^{-1}\right]_{q=\zeta}=1.$
5.6.
Here is the main result of this Section.
Theorem 5.8.
The power series $\mathsf{U}(h,\boldsymbol{a},q,z)$ given by (5.2) is well defined at $q$ given by unit roots $\zeta^{p^{s}}=1$. If $\zeta\neq 1$ is such unit root, then $\mathsf{U}(h,\boldsymbol{a},\zeta,z)$ is a Taylor series expansion of a rational function from $\Omega(z,\hbar,u_{1},\dots,u_{n})$ where
(5.13) $\hbar=\zeta^{h},\ \ u_{1}=\zeta^{a_{1}},\dots,u_{n}=\zeta^{a_{n}}$
In particular, these rational functions satisfy the congruences
(5.14) $\mathsf{U}(h,\boldsymbol{a},\zeta,z)=\mathsf{U}(h^{(s)},\boldsymbol{a}^{(s)},\zeta,z).$
where $\boldsymbol{a}^{(s)}=(a_{1}^{(s)},\dots,a^{(s)}_{n})$ and $h^{(s)}$ are natural numbers defined by $h^{(s)}=h\pmod{p^{s-1}}$, $a^{(s)}_{i}=a_{i}\pmod{p^{s-1}}$, $i=1,\dots,n$.
Proof.
Define $h^{(s)}\in\mathbb{N}$ by $h^{(s)}=h\pmod{p^{s-1}}$. We have $h=h^{(s)}+p^{s-1}h^{\prime}$ for some $h^{\prime}\in\mathbb{Z}_{p}$. By Theorem 3.1 we obtain:
$\widetilde{\Psi}(ph,p\boldsymbol{a},q,z)=\widetilde{\Psi}(ph^{(s)}+h^{\prime}p^{s},p\boldsymbol{a},q,z)=\left(\prod\limits_{i=h^{\prime}p^{s}}^{ph^{(s)}+h^{\prime}p^{s}-1}H(p\boldsymbol{a},i,q,z)\right)\widetilde{\Psi}(h^{\prime}p^{s},p\boldsymbol{a},q,z),$
where $H(p\boldsymbol{a},i,q,z)$ is a certain invertible matrix whose coefficients are rational functions of $z$ and (5.13). Similarly, we have
$\widetilde{\Psi}(h,\boldsymbol{a},q^{p},z^{p})=\left(\prod\limits_{i=h^{\prime}p^{s-1}}^{h^{(s)}+h^{\prime}p^{s-1}-1}H(\boldsymbol{a},i,q^{p},z^{p})\right)\widetilde{\Psi}(h_{1}p^{s-1},\boldsymbol{a},q^{p},z^{p})$
and therefore:
$\widetilde{\Psi}(ph,p\boldsymbol{a},q,z)\,\widetilde{\Psi}(h,\boldsymbol{a},q^{p},z^{p})^{-1}=$
$\left(\prod\limits_{i=h^{\prime}p^{s}}^{ph^{(s)}+h^{\prime}p^{s}-1}H(p\boldsymbol{a},i,q,z)\right)\widetilde{\Psi}(h^{\prime}p^{s},p\boldsymbol{a},q,z)\widetilde{\Psi}(h^{\prime}p^{s-1},\boldsymbol{a},q^{p},z^{p})^{-1}\left(\prod\limits_{i=h^{\prime}p^{s-1}}^{h^{(s)}+h^{\prime}p^{s-1}-1}H(\boldsymbol{a},i,q^{p},z^{p})\right)^{-1}$
By Theorem 5.7 at $q=\zeta$ the middle term in the last expression is trivial. We note that $H(p\boldsymbol{a},i,\zeta,z)=H(p\boldsymbol{a},i\pmod{p^{s}},\zeta,z)$ since it depends on $i$ via $\zeta^{i}=\zeta^{i\pmod{p^{s}}}$. Thus, we obtain:
$\left[\widetilde{\Psi}(ph,p\boldsymbol{a},q,z)\,\widetilde{\Psi}(h,\boldsymbol{a},q^{p},z^{p})^{-1}\right]_{q=\zeta}=\left(\prod\limits_{i=0}^{ph^{(s)}-1}H(p\boldsymbol{a},i,\zeta,z)\right)\left(\prod\limits_{i=0}^{h^{(s)}-1}H(\boldsymbol{a},i,\zeta^{p},z^{p})\right)^{-1}$
which is a product of finitely many matrices whose entries are rational functions in $z$ and $\zeta^{pa_{i}}$. We have $\zeta^{pa_{i}}=\zeta^{pa^{(s)}_{i}}$ where $a^{(s)}_{i}=a_{i}\mod p^{s-1}$. This finishes the proof. ∎
Corollary 5.9.
Let $\zeta\neq 1$ be a unit root satisfying $\zeta^{p^{s}}=1$, then we have
$\left.\mathsf{U}(h,\boldsymbol{a},q,z)\right|_{q=\zeta}=\left(\prod\limits_{i=0}^{ph^{(s)}-1}H(p\boldsymbol{a}^{(s)},i,\zeta,z)\right)\left(\prod\limits_{i=0}^{h^{(s)}-1}H(\boldsymbol{a}^{(s)},i,\zeta^{p},z^{p})\right)^{-1}$
5.7. Proof of Theorem 5.2
We are now ready to proceed to the proof of Theorem 5.2.
Proof.
By its definition (5.2), $\mathsf{U}(h,\boldsymbol{a},q,z)$ is map sending the fundamental solution $\Psi(h,\boldsymbol{a},q^{p},z^{p})$ to the fundamental solution $\Psi(ph,p\boldsymbol{a},q,z)$, i.e., it is an intertwiner between $q$-difference equations with these parameters. To show that it is a Frobenius intertwiner we need to check that $\mathsf{U}(h,\boldsymbol{a},q,z)\in\mathsf{GL}_{n}(\widehat{E}_{p})$. We have
$\prod\limits_{{\zeta\neq 1,}\atop{\zeta^{p^{s}}=1}}\,(q-\zeta)=[p^{s}]_{q}=\dfrac{1-q^{p^{s}}}{1-q}=\varphi_{p}(q)\varphi_{p}(q^{p})\dots\varphi_{p}(q^{p^{s-1}})$
where $\varphi_{p}(q)=1+q+\dots+q^{p-1}$ is the $p$ th cyclotomic polynomial. Theorem 5.8 therefore gives:
(5.15) $\mathsf{U}(h,\boldsymbol{a},q,z)=\mathsf{U}(h^{(l)},\boldsymbol{a}^{(l)},q,z)+[p^{l}]_{q}\,(\dots)$
where $\mathsf{U}(h^{(l)},\boldsymbol{a}^{(l)},q,z)\in\Omega(z)$ and $\dots$ stands for a power series in $z$ with coefficients regular at $q$ given by $p^{l}$-th unit roots. Since by our assummption $q=1+t$ with $|t|_{p}<1$ we have $[p^{l}]_{q}\equiv 0\pmod{p^{s}}$ for sufficiently large $l$, and therefore (5.15) implies that $\mathsf{U}(h,\boldsymbol{a},q,z)\pmod{p^{s}}$ is in $\Omega(z)$. As in Lemma 4.1, this implies that $\mathsf{U}(h,\boldsymbol{a},q,z)\in\mathsf{GL_{n}}(\hat{E}_{p})$. ∎
Remark.
In the scalar case case $n=1$ a similar result was recently obtained in [VM22].
6. Limit to cohomology
By Theorem 5.8 from the previous section at $q=\zeta$, where $\zeta\neq 1$ is a unit root satisfying $\zeta^{p^{s}}=1$, the Frobenius intertwiner $\mathsf{U}(h,\boldsymbol{a},q,z)$ specializes to a rational function of $z$. In this section we treat the remaining unit root $\zeta=1$.
For this, let us consider the sequence $q=1+p^{i},\ \ i=1,2,\dots$ By Theorem 5.2, for all $q$ in this sequence $U(h,\boldsymbol{a},q,z)$ is a Frobenius intertwiner. Moreover, since all these $q$ as rational $U(h,\boldsymbol{a},q,z)\in\mathsf{GL}_{n}(E_{p})$ where $E_{p}$ is the field of rational analytic elements, i.e., is the completion of $\mathbb{Q}(z)$ with respect to the Gauss norm. The limit $i\to\infty$ corresponds to the limit $q\to 1$ in the $p$-adic norm. In turn, it is well-known that in the limit $q\to 1$ the fundamental solution matrix $\Psi(h,\boldsymbol{a},q,z)$ of the quantum $q$-difference equation converges to the fundamental solution of the quantum differential equation of $X$, e.g., note that the $q$-hypergeometric functions (2.3) specialize to the standard hypergeometric series in this limit.
Thus, in the limit $q\to 1$ the intertwiner $U(h,\boldsymbol{a},q,z)$ converges to the Frobenius intertwiner for the hypergeometric differential equation. In this way we arrive at the classical results of Dwork [Dw69] on the Frobenius structures for the hypergeometric differential equations. In this Section we summarize the outcome of this limiting procedure.
6.1.
As $q\to 1$, the quantum difference equation of $X$ degenerates to the quantum differential equation, which governs the quantum cohomology of $X$. In this limit, the equivariant $K$-theory ring (3.1) is replaced by the equivariant cohomology ring
(6.1) $H^{\bullet}_{T}(X)=\mathbb{Z}[x,a_{1},\dots,a_{n},h]/(x-a_{1})\dots(x-a_{n}).$
and the gamma class (3.2) by:
(6.2) $\Gamma(h,\boldsymbol{a},x,z)=z^{x}\,\prod\limits_{i=1}^{n}\dfrac{\Gamma(x+h-a_{i})}{\Gamma(x+1-a_{i})\Gamma(h)}\in\hat{H}^{\bullet}_{T}(X).$
where $\Gamma$ denotes the classical Gamma function, and $\hat{H}^{\bullet}_{T}(X)$ is the appropriate completion of the cohomology ring containing the gamma functions. The normalized vertex function (3.3) (also known as $J$-function in quantum cohomology) now equals:
(6.3) $\widetilde{V}(h,\boldsymbol{a},x,z)=\sum\limits_{d=0}^{\infty}\,\Gamma(h,\boldsymbol{a},x+d,z)\,\in\,\hat{H}^{\bullet}_{T}(X)[[z]].$
A direct computation shows that this function satisfies the differential equation
(6.4) $P(\boldsymbol{a},h,z)\widetilde{V}(h,\boldsymbol{a},x,z)=(x-a_{1})\dots(x-a_{n})\Gamma(h,\boldsymbol{a},x,z)=0$
(the last equality is because of the relation in (6.1)), where
(6.5) $P(\boldsymbol{a},h,z)=z\prod\limits_{i=1}^{n}\,(D+h-a_{i})-\prod\limits_{i=1}^{n}(D-a_{i}),\ \ \ D=z\frac{d}{dz},$
is the hypergeometric differential operator. The components of the vertex function (6.3) in any basis of $H^{\bullet}_{T}(X)$ provide a basis of solutions for the hypergeometric differential equation (6.4). For example, the components of the vertex function (6.3) in the basis of $H^{*}_{T}(X)$ given by the $T$-fixed points $X^{T}$ are obtained by substituting $x=a_{i}$, $i=1,\dots,n$. In this way, we obtain a basis of solutions to (6.4) given by the classical hypergeometric series:
$\widetilde{V}(h,\boldsymbol{a},a_{i},z)=\Gamma(h,\boldsymbol{a},a_{i},z)\,\sum\limits_{d=0}^{\infty}\,\left(\prod\limits_{j=1}^{n}\dfrac{(a_{i}+h-a_{j})_{d}}{(a_{i}+1-a_{j})_{d}}\right)z^{d},\ \ i=1,\dots,n,$
where $(a)_{d}=a(a+1)\dots(a+d-1)$.
Let us denote $\mathsf{U}(h,\boldsymbol{a},z):=\left.\mathsf{U}(h,\boldsymbol{a},q,z)\right|_{q=1}$. By Theorem A.3 $\Gamma_{p,q}(n)\stackrel{{\scriptstyle q\to 1}}{{\longrightarrow}}\Gamma_{p}(n)$ and from Theorem 5.2 we obtain:
Theorem 6.1.
For any $\boldsymbol{a}=(a_{1},\dots,a_{n})\in\mathbb{Z}^{n}_{p},\ \ h\in\mathbb{Z}_{p}$ the power series $\mathsf{U}(h,\boldsymbol{a},z)$ takes values in $\mathsf{GL}_{n}(E_{p})$ and gives a Frobenius intertwiner from the hypergeometric system (6.5) with parameters $(a_{1},\dots,a_{n},h,z^{p})$ to the hypergeometric system with parameters $(pa_{1},\dots,pa_{n},ph,z)$. The constant term of this intertwiner is the operator acting in $\hat{H}^{\bullet}_{T}(X)$ as an operator of multiplication by the cohomology class class
(6.6) $\mathsf{U}(h,\boldsymbol{a},0)=\prod\limits_{j=1}^{n}\,\dfrac{\Gamma_{p}(px+ph-pa_{j})}{\Gamma_{p}(px-pa_{j})\Gamma_{p}(ph)},$
where $\Gamma_{p}$ is the Morita gamma function.
The $q\to 1$ limit of Corollary 5.4 also gives:
Corollary 6.2.
If $a_{1},\dots,a_{n}$ are pairwise distinct, then the constant term of the Frobenius intertwiner $\mathsf{U}(h,\boldsymbol{a},z)$ is diagonal in the basis of $H^{*}_{T}(X)$ given by the classes of $T$-fixed points with the eigenvalues
(6.7) $\mathsf{U}(h,\boldsymbol{a},0)_{i,i}=\prod\limits_{j=1}^{n}\,\dfrac{\Gamma_{p}(p(a_{i}-a_{j}+h))}{\Gamma_{p}(p(a_{i}-a_{j}))\Gamma_{p}(ph)},\ \ \ i=1,\dots,n.$
In this way we arrive at the description of the Frobenius intertwines for the $p$-adic hypergeometric equations discovered by Dwork [Dw69, Dw89] see also [Ked21] where this result was recently revisited by other tools.
Recall that for $n\in p\mathbb{Z}_{p}$ the Morita gamma function satisfies the identities $\Gamma_{p}(n)=-\Gamma_{p}(1+n)=1/\Gamma_{p}(-n)$. Using this, up to a constant scalar multiple independent of $x$, (6.6) can we written in the form
$\mathsf{U}(h,\boldsymbol{a},0)=\prod\limits_{i=1}^{n}\,\Gamma_{p}(px+ph-pa_{j})\Gamma_{p}(pa_{j}-px)$
The arguments of the gamma functions in this product are exactly the $p$-multiples of the Chern roots of the tangent bundle $TX$, i.e., they are the Chern roots of $TX^{(p)}:=\psi^{p}(TX)$ where $\psi^{p}$ is the $p$-th Adams operation. We thus can write this formally as
Proposition 6.3.
The constant term of the intertwiner $\mathsf{U}(h,\boldsymbol{a},{\boldsymbol{z}})$ acts in the equivariant cohomology $H^{*}_{T}(X)$ as multiplication by the cohomology class
(6.8) $\mathsf{U}(h,\boldsymbol{a},0)=\Gamma_{p}(TX^{(p)})$
Here, the cohomology class $\Gamma_{p}({\mathcal{V}})$ of a vector bundle $\mathcal{V}$ is defined as in (5.7) with $\Gamma_{p,q}$ replaced by the Morita gamma $\Gamma_{p}$. We expect that (6.8) holds for the constant term of the Frobenius intertwiner of the quantum differential equations associated with more general varieties than $X$. The last Proposition is the cohomological limit of Proposition 5.5.
7. Limit to $X=\mathbb{P}^{n-1}$: Frobenius structure for generalized Bessel equations
7.1.
The equivariant quantum differential equation and its solutions for $X=\mathbb{P}^{n-1}$ arise as the limit $\hbar\to\infty$ of the corresponding equations and solutions for $T^{*}\mathbb{P}^{n-1}$ (we recall that the equivariant parameter $\hbar$ is the character of $T$ corresponding to the dilation of the cotangent directions of $T^{*}\mathbb{P}^{n-1}$ under the $T$-action) [SV24]. Thus, the formal limit $\hbar\to\infty$ of our results in the previous section gives the Frobenius intertwiners for the $T$-equivariant quantum differential equations of projective spaces.
Further, in the non-equivariant limit, when we ignore the $T$-action on $X=\mathbb{P}^{n-1}$, the quantum differential equation is nothing but the generalized Bessel equation (and the classical Bessel differential equation for $n=2$). The non-equivariant limit corresponds to the case $\boldsymbol{a}=(0,\dots,0)$. Specializing at this point, we arrive at the Frobenius structure for Bessel equations studied in [Dw74, Sp80]. This section is a summary of these limits.
7.2.
As before, let $T=(\mathbb{C}^{\times})^{n}$ be the torus acting on $\mathbb{C}^{n}$ by scaling the coordinate lines with characters $a_{1},\dots,a_{n}$. We consider the induced action on $X=\mathbb{P}^{n-1}=\mathbb{P}(\mathbb{C}^{n})$. The equivariant cohomology ring has the form
(7.1) $H^{\bullet}_{T}(X)=\mathbb{Z}[x,a_{1},\dots,a_{n}]/(x-a_{1})\dots(x-a_{n}).$
Taking the formal limit $\hbar\to\infty$ of (6.2) we obtain the gamma class for $X=\mathbb{P}^{n-1}$:
$\Gamma(x,\boldsymbol{a},z)=\dfrac{z^{x}}{\prod\limits_{i=1}^{n}\Gamma(x+1-a_{i})}\in\hat{H}^{\bullet}_{T}(X).$
The normalized equivariant cohomological vertex function (the $J$-function) of $\mathbb{P}^{n-1}$ equals:
$\widetilde{V}(x,\boldsymbol{a},z)=\sum\limits_{d=0}^{\infty}\,\Gamma(x+d,\boldsymbol{a},z)\,\in\,\hat{H}^{\bullet}_{T}(X)[[z]].$
More explicitly
$\widetilde{V}(x,\boldsymbol{a},z)=\dfrac{z^{x}}{\prod\limits_{i=1}^{n}\Gamma(x+1-a_{i})}\,\sum\limits_{d=0}^{\infty}\,\dfrac{z^{d}}{(x-a_{1}+1)_{d}\dots(x-a_{n}+1)_{d}}.$
This power series satisfies the differential equation
$P(\boldsymbol{a},z)\widetilde{V}(x,\boldsymbol{a},z)=(x-a_{1})\dots(x-a_{n})\Gamma(x,\boldsymbol{a},z),$
where $P(\boldsymbol{a},z)$ is the differential operator
(7.2) $P(\boldsymbol{a},z)=z-(D-a_{1})\cdots(D-a_{n}),\ \ \ D=z\dfrac{d}{dz}$
and therefore in the ring (7.1) we have
(7.3) $P(\boldsymbol{a},z)\widetilde{V}(x,\boldsymbol{a},z)=0.$
Let $\mathsf{U}(h,\boldsymbol{a},z)$ be as in Theorem 6.1 and let $\mathsf{U}(\boldsymbol{a},z)$ denotes its limit as $\hbar\to\infty$. Then we obtain
Theorem 7.1.
[…]
(7.4) $\mathsf{U}(\boldsymbol{a},0)=\prod\limits_{j=1}^{n}{\Gamma_{p}}(px-pa_{j}).$
We note that the arguments of the gamma functions in the product (7.4) are the $p$ multiples of the Chern roots of the tangent bundle $TX$. Thus, the universal formula (6.8) still holds.
7.3.
At $\boldsymbol{a}=(0,\dots,0)$ the equivariant cohomology of $X=\mathbb{P}^{n-1}$ specializes at the usual (non-equivariant) cohomology:
(7.5) $H^{\bullet}(X)=\mathbb{Z}[x]/(x)^{n}.$
The gamma class and the vertex function have the form
$\Gamma(x,z)=\dfrac{z^{x}}{\Gamma(x+1)^{n}},$
and
(7.6) $\widetilde{V}(x,z)=\sum\limits_{d=0}^{\infty}\,\Gamma(x+d,z)=\dfrac{z^{x}}{\Gamma(x+1)^{n}}\,\sum\limits_{d=0}^{\infty}\,\dfrac{z^{d}}{(x+1)^{n}_{d}},$
respectively. The vertex function solves
$P(z)\widetilde{V}(x,z)=x^{n}\,\Gamma(x,z)=0$
where
(7.7) $P(z)=z-D^{n}$
The differential equation defined by (7.7) is the generalized Bessel equation. For $n=2$ this is the classical Bessel equation. The components of $\widetilde{V}(x,z)$ in any basis of cohomology (7.5) provide a basis of solutions for these equations. In this case, the only natural basis is given by the powers of the hyperplane class $x$:
(7.8) basis $\textrm{basis of}\ H^{\bullet}(X)=\{1,x,x^{2},\dots,x^{n-1}\}.$
In order to compute the corresponding components it is enough to expand (7.6) in the Taylor expansion up to degree $x^{n}$:
$\widetilde{V}(x,z)=\widetilde{V}_{0}(z)+\widetilde{V}_{1}(z)x+\dots+\widetilde{V}_{n-1}x^{n-1}$
The functions $\widetilde{V}_{0}(z),\dots,\widetilde{V}_{n}(z)$ provide a basis of solutions to the generalized Bessel equation. Note that (7.6) includes a multiple $z^{x}$ which has the following expansion:
$z^{x}=\sum\limits_{m=0}^{n-1}\,\dfrac{\ln(z)^{m}}{m!}x^{m}$
and thus the solutions may be written in the form
(7.14) $\ \ \ \ \ \ \ \ \tiny(\widetilde{V}_{0}(z),\dots,\widetilde{V}_{n-1}(z))=({V}_{0}(z),\dots,{V}_{n-1}(z))\cdot\left(\begin{array}[]{ccccc}1&\ln(z)&\ln(z)^{2}/2&..&\ln(z)^{n-1}/(n-1)!\\ 0&1&\ln(z)&..&\ln(z)^{n-2}/(n-2)!\\ 0&0&1&..&\dots\\ ..&..&..&..&\dots\\ 0&0&0&..&1\end{array}\right)$
where $V_{i}(z)$ are some power series in $z$. In particular, $\tilde{V}_{0}(z)$ does not contain logarithmic terms. For $n=2$, $\tilde{V}_{0}(z)$ is the power series representing the classical Bessel function of the first kind.
7.4.
When all $a_{i}=0$, the Theorem 7.3 specializes to:
Theorem 7.2.
The generalized Bessel equation admits Frobenius intertwiner $z\to z^{p}$ with the constant term acting in $H^{\bullet}(X)$ as the operator of multiplication by the cohomology class
(7.15) $\mathsf{U}(0)={\Gamma_{p}}(px)^{n}$
By construction, the intertwiner of Theorem 7.3 maps the solutions $\tilde{V}(x,z^{p})$ to the solutions $\tilde{V}(px,z)$. Let $p^{\mathrm{deg}}$ be the operator of the cohomological grading, acting on the cohomology by $p^{\mathrm{deg}}(x^{k})=p^{k}x^{k}$. The above theorem says that
${\Gamma_{p}}(x)^{n}\cdot p^{\mathrm{deg}}$
is the zero term of the intertwiner between $\tilde{V}(x,z^{p})$ and $\tilde{V}(x,z)$ (note that as in (7.14), the operators act by multiplication from the right in our conventions). In other words this gives an automorphism of the corresponding differential equation. We thus conclude with
Theorem 7.3.
The generalized Bessel equation admits unique Frobenius automorphism acting by $z\to z^{p}$. The constant term of the automorphsim acts in cohomology as the operator
(7.16) $\mathsf{U}(0)={\Gamma_{p}}(x)^{n}\cdot p^{\mathrm{deg}}.$
Let us compute the matrix (7.16) explicitly in the basis (7.8). If $\Psi(x,z)=\Psi_{0}(z)+\Psi_{1}(z)x+\dots+\Psi_{n-1}x^{n-1}$, then the convention is that the matrices of the operators act on the row vector $(\Psi_{0}(z),\Psi_{1}(z),\dots,\Psi_{n-1})$ by multiplication from the right. For instance, multiplication by $x$ acts by $x\Psi(x,z)=\Psi_{0}(z)x+\Psi_{1}(z)x^{2}+\dots+\Psi_{n-2}x^{n-1}$, and thus the corresponding matrix is
$x=\left(\begin{array}[]{cccccc}0&1&0&0&\dots&0\\ 0&0&1&0&\dots&0\\ 0&0&0&1&\dots&0\\ ..&..&..&..&\dots&1\\ 0&0&0&0&\dots&0\\ \end{array}\right).$
Let us recall that the logarithm of the Morita gamma function has the following Taylor expansion, see for instance Sections 58 and 61 in [Sh]:
$\log\Gamma_{p}(x)=\Gamma_{p}^{\prime}(0)x-\sum\limits_{m\geqslant 2}\,\dfrac{\zeta_{p}(m)}{m}\,x^{m}$
where $\zeta_{p}(m)$ denote the values of $p$-adic zeta function. It is known that $\zeta_{p}(m)=0$ for even $m$, so only the odd powers of $x$ appear in the expansion. Thus, in our conventions we obtain that the operator of multiplication by $\log\Gamma_{p}(x)^{n}$ is given by the $n\times n$ matrix
$\log\Gamma_{p}(x)^{n}=\left(\begin{array}[]{cccccc}0&\nu_{1}&0&\nu_{3}&\dots&\dots\\ 0&0&\nu_{1}&0&\nu_{3}&\dots\\ 0&0&0&\nu_{1}&\dots&\dots\\ ..&..&..&..&\dots&0\\ ..&..&..&..&\dots&\nu_{1}\\ 0&0&0&0&\dots&0\\ \end{array}\right),$
where we denoted $\nu_{1}=n\Gamma_{p}^{\prime}(0)$ and $\nu_{m}=-n\,\zeta_{p}(m)/m$ for $m>2$. We also obviously have
$p^{\mathrm{deg}}=\left(\begin{array}[]{cccccc}1&0&0&0&\dots&0\\ 0&p&0&0&\dots&0\\ 0&0&p^{2}&0&\dots&0\\ ..&..&..&..&\dots&0\\ ..&..&..&..&p^{n-2}&0\\ 0&0&0&0&\dots&p^{n-1}\\ \end{array}\right).$
Finally, we obtain that the constant term of the automorphism has the following matrix:
$\mathsf{U}(0)={\Gamma_{p}}(x)^{n}\cdot p^{\mathrm{deg}}=\exp\left(\begin{array}[]{cccccc}0&\nu_{1}&0&\nu_{3}&\dots&\dots\\ 0&0&\nu_{1}&0&\nu_{3}&\dots\\ 0&0&0&\nu_{1}&\dots&\dots\\ ..&..&..&..&\dots&0\\ ..&..&..&..&\dots&\nu_{1}\\ 0&0&0&0&\dots&0\\ \end{array}\right)\cdot\left(\begin{array}[]{cccccc}1&0&0&0&\dots&0\\ 0&p&0&0&\dots&0\\ 0&0&p^{2}&0&\dots&0\\ ..&..&..&..&\dots&0\\ ..&..&..&..&p^{n-2}&0\\ 0&0&0&0&\dots&p^{n-1}\\ \end{array}\right)$
and the exponent of a nilpotent operator is, of course, well defined.
Example.
For the case of the classical Bessel equation, corresponding to $n=2$, this matrix has the form
(7.19) $\mathsf{U}(0)=\left(\begin{array}[]{ll}1,&p\,\Gamma^{\prime}_{p}(0)\\ 0,&p\end{array}\right).$
The existence of the Frobenius automorphism for the Bessel equation was discovered by Dwork [Dw74]. In particular, in Lemma 3.2 of [Dw74] he gave an explicit formula for the matrix $\mathsf{U}(0)$ in terms of certain converging $p$-adic series. One checks that Dwork’s matrix $\mathfrak{U}(0)$ coincides with (7.19) after transposition. In particular, his series for the coefficient $\mathfrak{U}_{2,1}(0)=\gamma$ converges to $p\Gamma^{\prime}_{p}(0)$.
The action of the Frobenius automorphism for the generalized Bessel equations ($n>2$) was further studied by Sperber in [Sp80]. The results of this Section are also in full agreement with these results.
We note that the values of $p$-adic zeta function appear naturally in our analysis of hypergeometric equations when we pass from the equivariant cohomology to the non-equivariant cohomology. Namely, the values of $p$-adic zeta functions appear as the coefficients of the expansion of the cohomology class $\Gamma_{p}(TX^{(p)})$ is some basis of $H^{*}(X)$. We expect the values of $p$-adic zeta appearing in the study of the mirror symmetry for more general examples of $X$, e.g., [BV23], should be treated in the same way.
8. Numerical examples of Theorem 5.2
8.1.
Let ${\Psi}(h,\boldsymbol{a},q,z)$ be the fundamental solution matrix (2.11). For simplicity we assume that all parameters are rational. As in (5.1) we look for the Frobenius intertwiner matrix in the form
(8.1) $\mathsf{U}(z)={\Psi}(ph,p\boldsymbol{a},q,z)\Lambda{\Psi}(h,\boldsymbol{a},q^{p},z^{p})^{-1}.$
where $\Lambda$ is a constant matrix. In order to be a Frobenius intertwiner, we need to find $\Lambda$ for which (8.1) is an element of $\mathsf{GL}_{n}(E_{p})$. Lemma 4.1 says that $\mathsf{U}(z)$ must be a power series
$\mathsf{U}(z)=\mathsf{U}_{0}+\mathsf{U}_{1}z+\dots\in\mathsf{Mat}_{n}(\mathbb{Q})[[z]]$
such that for any natural number $s$, there exists a natural number $m_{s}$ for whichLemma 4.1 says that $\mathsf{U}(z)\pmod{p^{s}}$ must be a rational function. Denominators of these rational functions are controlled by the singularities of the corresponding differential equations. In our case, the qde has a singularity at $z=1$ and therefore the denominators must be powers of $(1-z)$.
(8.2) $(1-z)^{m_{s}}\,\mathsf{U}(z)\pmod{p^{s}}\in\mathsf{Mat}_{n}(\mathbb{Z})[z].$
In words: after multiplication by sufficiently large power of $(1-z)$ the matrix elements of $\mathsf{U}(z)$ reduced $\pmod{p^{s}}$ truncate to polynomials in $z$. This condition is very strong and can be used to effectively discover non-trivial Frobenius intertwiners as we are about to show.
8.2.
Note that if the parameters $\boldsymbol{a}=(a_{1},\dots,a_{n})$ are pairwise distinct the matrix $\Lambda$ must be diagonal. Indeed, the fundamental solution matrix ${\Psi}(\boldsymbol{a},\hbar,q,z)$ contains logarithmic terms given by matrix $\mathsf{E}(z)$ in (2.12). The intertwiner (8.1) is a power series in $z$ only if these log terms are canceled. It is elementary to check that for the pairwise distinct $a_{i}$ this happens only if $\Lambda$ is diagonal. Since $\Lambda$ is defined up to a scalar factor, we may assume further that $\Lambda_{n,n}=1$ so
(8.8) $\Lambda=\left(\begin{array}[]{ccccc}c_{1}&0&0&..&0\\ 0&c_{2}&0&..&0\\ ..&..&..&..&..\\ 0&0&..&c_{n-1}&0\\ 0&0&0&..&1\end{array}\right)$
for some unknown $c_{1},\dots,c_{n-1}\in\mathbb{Z}_{p}$.
8.3.
Let us assume $n=2$. In this case
(8.11) $\Lambda=\left(\begin{array}[]{cc}c&0\\ 0&1\end{array}\right).$
Let us pick a prime $p=3$, set $q=1+p=4$, and pick some random values of $a_{1},a_{2},h\in\mathbb{Z}_{p}$, for instance
(8.12) $h=\dfrac{1}{2},\ \ a_{1}=\dfrac{1}{5},\ \ a_{2}=0.$
(Since the solutions only depend on the difference $a_{1}-a_{2}$ we can always assume $a_{2}=0$). For our computer check we use integral approximations of these numbers say up to $p^{7}$:
$\bar{h}=h\pmod{3^{7}}=1094,\ \ \bar{a_{1}}=a_{1}\pmod{3^{7}}=875,\ \ \ \bar{a_{2}}=0.$
We now determine the constant $c\in\mathbb{Z}_{p}$ in (8.11) from the rationality condition (8.2) in a chain of steps. For this we assume that $c$ has the following $p$-adic expansion:
$c=c^{(0)}+c^{(1)}3+c^{(2)}3^{2}+\dots$
where $c^{(i)}\in\{0,1,2\}$.
Step 1: determination of $c^{(0)}$ and $c^{(1)}.$
Using a computer we approximate the power series ${\Psi}(\boldsymbol{a},\hbar,q,z)$ by series expansion (2.11) up to $O(z^{20})$. Substituting this approximation to (8.1) and trying the integers $0,1,2$ for the values of $c^{(0)}$ and $c^{(1)}$ we find that $\mathsf{U}(z)\pmod{p}$ exists (i.e. no powers of $p$ in the denominators) only if $c^{(0)}=1$ and $c^{(1)}=0$, in which case we find that
(8.15) $\mathsf{U}(z)\pmod{p}=\left(\begin{array}[]{cc}1+O\left({z}^{20}\right)&O\left({z}^{20}\right)\\ \par\vskip 6.0pt plus 2.0pt minus 2.0pt\cr O\left({z}^{20}\right)&1+O\left({z}^{20}\right)\end{array}\right)$
We note that $\mathsf{U}(z)\pmod{p}$ given by (8.15) does not depend on the value of $c^{(2)}$ and higher.
Step 2: determination of $c^{(2)}.$
If $c^{(2)}=0$, using a computer we find
$\mathsf{U}(z)\pmod{p}=\left[\begin{array}[]{cc}1+O\left({z}^{20}\right)&O\left({z}^{20}\right)\\ \par\vskip 6.0pt plus 2.0pt minus 2.0pt\cr{z}^{12}+2\,{z}^{15}+O\left({z}^{20}\right)&1+O\left({z}^{20}\right)\end{array}\right]$
If $c^{(2)}=1$ we find
$\mathsf{U}(z)\pmod{p}=\left[\begin{array}[]{cc}1+O\left({z}^{20}\right)&O\left({z}^{20}\right)\\ \par\vskip 6.0pt plus 2.0pt minus 2.0pt\cr O\left({z}^{20}\right)&1+O\left({z}^{20}\right)\end{array}\right]$
If $c^{(2)}=2$ we find
$\mathsf{U}(z)\pmod{p}=\left[\begin{array}[]{cc}1+O\left({z}^{20}\right)&O\left({z}^{20}\right)\\ \par\vskip 6.0pt plus 2.0pt minus 2.0pt\cr 2\,{z}^{12}+{z}^{15}+O\left({z}^{20}\right)&1+O\left({z}^{20}\right)\end{array}\right]$
We see that the only value of $c^{(2)}$ for which the result is a polynomial in $z$ is $c^{(2)}=1$. These results do not depend on the value of the coefficient $c^{(3)}$ and higher which are to be determined in the next step.
Step 3: determination of $c^{(3)}.$
If $c^{(3)}=0$ we find:
$(1-z)^{2}\mathsf{U}(z)\!\!\pmod{p^{2}}=\left[\begin{array}[]{cc}1+z+{z}^{2}+O\left({z}^{20}\right)&O\left({z}^{20}\right)\\ \par\vskip 6.0pt plus 2.0pt minus 2.0pt\cr 6\,z+3\,{z}^{12}+3\,{z}^{13}+3\,{z}^{14}+6\,{z}^{15}+6\,{z}^{16}+6\,{z}^{17}+O\left({z}^{20}\right)&1+7\,z+{z}^{2}+O\left({z}^{20}\right)\end{array}\right]$
If $c^{(3)}=1$ we find:
$(1-z)^{2}\mathsf{U}(z)\!\!\pmod{p^{2}}=\left[\begin{array}[]{cc}1+z+{z}^{2}+O\left({z}^{20}\right)&O\left({z}^{20}\right)\\ \par\vskip 6.0pt plus 2.0pt minus 2.0pt\cr 6\,z+O\left({z}^{20}\right)&1+7\,z+{z}^{2}+O\left({z}^{20}\right)\end{array}\right]$
If $c^{(3)}=2$ we find:
$(1-z)^{2}\mathsf{U}(z)\!\!\pmod{p^{2}}=\left[\begin{array}[]{cc}1+z+{z}^{2}+O\left({z}^{20}\right)&O\left({z}^{20}\right)\\ \par\vskip 6.0pt plus 2.0pt minus 2.0pt\cr 6\,z+6\,{z}^{12}+6\,{z}^{13}+6\,{z}^{14}+3\,{z}^{15}+3\,{z}^{16}+3\,{z}^{17}+O\left({z}^{20}\right)&1+7\,z+{z}^{2}+O\left({z}^{20}\right)\end{array}\right]$
We conclude that there is unique value of $c^{(3)}=1$ for which the result is a polynomial in $z$.
Continuing this process we will find that at $m$-th step there is always unique value of $c^{(m)}$ for which rationality condition (8.2) is met. Up to the order $9$ these method gives the following value
[…]
8.4.
Theorem 5.2 and Corollary 5.4 give the following formula for the constant $c$:
$c=\dfrac{\Lambda_{1,1}}{\Lambda_{2,2}}=\dfrac{\Gamma_{p,q}(p(a_{1}-a_{2}+h))\Gamma_{p,q}(p(a_{2}-a_{1}))}{\Gamma_{p,q}(p(a_{2}-a_{1}+h))\Gamma_{p,q}(p(a_{2}-a_{1}))}$
at (8.12) we have
$c=\dfrac{\Gamma_{p,q}(\frac{21}{10})\Gamma_{p,q}(-\frac{3}{5})}{\Gamma_{p,q}(\frac{9}{10})\Gamma_{p,q}(\frac{3}{5})}.$
For $p=3$ and $q=1+p=4$ we compute:
[…]
in full agreement with the value obtained from the numerical experiment (8.16).
The author have performed the same numerical experiments for numerous other rational values of parameters $a_{1},a_{2}$ and $h$ in $\mathbb{Z}_{p}$. In all these cases the results are in full agreement with Theorem 5.2. Same method can be used for ranks $n>2$, in which case we have to determine the numerical values of $n-1$ constants in (8.8). Again, we find full agreement with Theorem 5.2.
9. Concluding remarks and future directions
9.1.
We expect that the Frobenius intertwiner can be defined enumeratively as a partition function counting the quasimaps from $\mathbb{P}^{1}$ to a Nakajima variety $X$. For this, one can consider the quasimaps with two different relative conditions at $0\in\mathbb{P}^{1}$ and $\infty\in\mathbb{P}^{1}$. The corresponding moduli space of quasimaps is represented by Fig.1, we refer to [Oko17] and also [PSZ] where the pictorial notations for various quasimap moduli spaces were introduced.
Figure 1. Frobenius intertwiner as a partition function of relative quasimaps
At $0\in\mathbb{P}^{1}$ we have the usual relative condition on quasimaps as defined in [Oko17]. We recall that the relative boundary condition at $0\in\mathbb{P}^{1}$ means that $\mathbb{P}^{1}$ is allowed to deform into a chain of rational “bubbles” with simple nodal intersections. This chain of bubbles is represented by the blue color in Fig. 1.
It is known, Theorem 8.1.16 in [Oko17], that the partition function of relative quasimaps from the chain of bubbles at $0\in\mathbb{P}^{1}$ is a fundamental solution matrix $\Psi(\hbar,{\boldsymbol{u}},q,z)$ of the quantum difference equation associated with $X$, normalized so that
$\Psi(\hbar,{\boldsymbol{u}},q,z)=1+O(z)\in K_{T}(X)^{\otimes 2}[[z]],$
in terminology of [Oko17] $\Psi({\boldsymbol{u}},\hbar,q,z)$ is also referred to as the capping operator.
At $\infty\in\mathbb{P}^{1}$ we also impose a relative condition but require, in addition, the relative quasimaps to be $\mathbb{Z}/p\mathbb{Z}$-invariant. More precisely, recall that the moduli space of relative quasimaps is defined as a stack quotient of quasimaps from a chain of $\mathbb{P}^{1}$’s by the automorphism group (×$(\mathbb{C}^{\times})^{l}$ where $l$ is the length of the chain, see Section 6.2 in [Oko17]. The moduli space of relative quasimaps at $\infty\in\mathbb{P}^{1}$ is constructed as follows: first we consider the space of quasimaps from a chain of rational curves of length $l$ which are invariant with respect to $\mathbb{Z}/p\mathbb{Z}$, which acts in each “bubble” (the component of the chain) via multiplication by $p$ th roots of unity. Then we define the $\mathbb{Z}/p\mathbb{Z}$-invariant quasimaps as the stack quotient of this moduli space by (×$(\mathbb{C}^{\times})^{l}$. This “$\mathbb{Z}/p\mathbb{Z}$-invariant chain of bubbles” is represented by the red color in Fig. 1.
By the $\mathbb{Z}/p\mathbb{Z}$-invariance, the degrees of the quasimaps allowed in this picture are all multiples of $p$. In particular, the corresponding partition function is a power series in $z^{p}$. For appropriately defined moduli space, the corresponding partition function arising at $\infty\in\mathbb{P}^{1}$ is equal to $\Psi({\boldsymbol{u}},\hbar,q^{p},z^{p})^{-1}$. The inverse arises due to transposition and inversion of character $q\to 1/q$ at $\infty\in\mathbb{P}^{1}$.
Finally, the contribution of degree 00 - quasimaps which correspond to the “constant” quasimaps from the parameterized component $\mathbb{P}^{1}$ can be computed by $\mathbb{C}^{\times}_{q}$ equivariant localization similarly to how it was done for the shift operators in Section 8.2 of [Oko17], in particular see Lemma 8.2.12. The resulting contribution has the form
$\dfrac{\Phi(ph,p\boldsymbol{a},px,q,z)}{\Phi(h,\boldsymbol{a},x,q^{p},z^{p})}\sim\Gamma_{p,q}\Big{(}(q^{p}-\hbar^{p})T^{1/2}X^{(p)}\Big{)},$
where $T^{1/2}X$ denotes the corresponding choice of the polarization bundle of $X$. The numerator of this fraction is the localization contribution at $0\in\mathbb{P}^{1}$ and the denominator is at $\infty\in\mathbb{P}^{1}$. Combining all this together we obtain that the corresponding partition function is
$\Psi({\boldsymbol{u}},\hbar,q,z)\Gamma_{p,q}\Big{(}(q^{p}-\hbar^{p})T^{1/2}X^{(p)}\Big{)}\Psi({\boldsymbol{u}},\hbar,q^{p},z^{p})^{-1}$
which is exactly the Frobenius intertwiner (5.2).
9.2.
The idea of $\mathbb{Z}/p\mathbb{Z}$ invariant quasimap moduli is similar to the one used in [HL24] to investigate the quantum Steenrod operators in quantum cohomology. It was shown in [HL24] that for a large class of varieties $X$, which includes the Nakajima varieties, the quantum Steenrod operations coincide with the $p$-curvature of the quantum connection.
The $q$-difference connections also have well defined $p$-curvature. Namely, the analog of $p$-curvature of $q$-difference connection (2.10) is:
$\mathbf{C}^{(p)}(\zeta,z)=\left.\Big{(}\mathbf{M}(zq^{p-1})\dots\mathbf{M}(zq)\mathbf{M}(z)\Big{)}\right|_{q=\zeta}$
where $\zeta$ is a non-trivial unit root of order $p$. In [KS] we showed that the $p$-curvature of the quantum connection of $X$ arises as the first non-trivial term in the $p$-adic expansion of $\mathbf{C}^{(p)}(\zeta,z)$.
Let us show that $p$-curvature $\mathbf{C}^{(p)}(\zeta,z)$ is similar to $\mathbf{M}(z^{p})$. Indeed, iterating the $q$-difference equation (2.10) $p$-times we obtain:
$\Psi(q,zq^{p})=\mathbf{M}(zq^{p-1})\dots\mathbf{M}(z)\Psi(z)$
also, rising all variables in (2.10) to their $p$-th powers we obtain:
$\Psi(q^{p},z^{p}q^{p})=\mathbf{M}(z^{p})\Psi(q^{p},z^{p})$
Dividing the first equation by the second, and specializing at $q=\zeta$ we arrive at:
$\mathsf{U}(\zeta,z)\mathbf{M}(z^{p})\mathsf{U}(\zeta,z)^{-1}=\mathbf{C}^{(p)}(\zeta,z).$
In particular, the spectrum of the $q$-deformed $p$-curvature $\mathbf{C}^{(p)}(\zeta,z)$ coincides with the spectrum of $\psi^{p}$-twisted quantum multiplication $\mathbf{M}(z^{p})$. This result is a $q$-version of the isospectrality recently proven in [EV24] in the differential case using different methods.
Acknowledgements
I thank S. Bai, P. Etingof, V. Golyshev, K. Kedlaya, J. Hee Lee, A. Oblomkov, A. Petrov, D. Vargas-Montoya, M. Vlasenko and especially A. Okounkov for very inspiring discussions. I thank A. Varchenko for introducing me to the subject. This note is the extended version of the talks [Sm23] and [Sm24], I thank the organizers for these opportunities. This work is supported by NSF grants DMS - 2054527 and DMS-2401380.
Appendix A $p$-adic $q$-deformations
In this section we recall standard notations of $q$-calculus. We also obtain results about $p$-adic $q$-gamma function which we we need in this paper.
A.1. $q$-numbers
For $n\in\mathbb{N}$ we define the $q$-number by:
$[n]_{q}=\dfrac{1-q^{n}}{1-q}=1+q+\dots+q^{n-1}.$
Let $[n]_{q}!=[1]_{q}[2]_{q}\dots[n]_{q}$ denote the $q$-factorial and
(A.1) $\binom{n}{k}_{\!\!q}=\dfrac{[n]_{q}!}{[k]_{q}![n-k]_{q}!}$
Let us define the reciprocal of the $q$-gamma functions as formal product:
$(x,q)_{\infty}=\prod\limits_{i=0}^{\infty}(1-xq^{i})$
We also define finite products
$(x,q)_{d}=\dfrac{(x,q)_{\infty}}{(xq^{d},q)_{\infty}}=(1-x)(1-xq)\dots(1-xq^{d-1})$
and
(A.2) $[x,q]_{d}=(1-qx)\dots(1-q^{d-1}x)=\dfrac{(xq,q)_{\infty}}{(xq^{d},q)_{\infty}}$
A.2. $q$-hypergeometric series at unit roots
Let us recall the $q$-binomial theorem:
$\dfrac{(z\hbar,q)_{\infty}}{(z,q)_{\infty}}=\sum\limits_{d=0}^{\infty}\,\dfrac{(\hbar,q)_{d}}{(q,q)_{d}}z^{d}$
At $\hbar=q^{l}$, $l\in\mathbb{N}$ it is convenient to write the coefficients as
(A.3) $\dfrac{(\hbar,q)_{d}}{(q,q)_{d}}=\dfrac{\frac{(\hbar,q)_{\infty}}{(\hbar q^{d},q)_{\infty}}}{\frac{(q,q)_{\infty}}{(qq^{d},q)_{\infty}}}=\dfrac{(\hbar,q)_{\infty}}{(q,q)_{\infty}}\dfrac{(qq^{d},q)_{\infty}}{(\hbar q^{d},q)_{\infty}}$
Using notation (A.2) at $\hbar=q^{l}$ we thus write:
(A.4) $\dfrac{(q^{l},q)_{d}}{(q,q)_{d}}=\dfrac{[q^{d},q]_{l}}{[1,q]_{l}}$
At the same time at $\hbar=q^{l}$ we have
$\dfrac{(z\hbar,q)_{\infty}}{(z,q)_{\infty}}=\dfrac{1}{(z,q)_{l}}$
Combining all this together, we arrive at the following version of the $q$-binomial theorem:
(A.5) $\dfrac{1}{(z,q)_{l}}=\sum\limits_{d=0}^{\infty}\,\dfrac{[q^{d},q]_{l}}{[1,q]_{l}}z^{d}$
Note that the coefficients of the power series (A.5) are polynomials in $q$ and in particular, can be specialized at unit roots. Let $p$ be a prime and let $\zeta\neq 1$ be a unit root $\zeta^{p^{s}}=1$ for some natural number $s$. We may assume that the order of $\zeta$ is $p^{m}$ with $1\leqslant m\leqslant s$. Let $h^{\prime}$ be a natural number. Then, applying (A.5) for $l=h^{\prime}p^{s}$ we have
$\left.\dfrac{1}{(z,q)_{h^{\prime}p^{s}}}\right|_{q=\zeta}=\dfrac{1}{(1-z^{p^{m}})^{h^{\prime}p^{s-m}}}=\sum\limits_{i=0}^{\infty}\,\left.\dfrac{[q^{i},q]_{h^{\prime}p^{s}}}{[1,q]_{h^{\prime}p^{s}}}\right|_{q=\zeta}z^{i}$
At the same time, since $\zeta^{p}$ is a primitive unit roots of order $p^{m-1}$ applying (A.5) for $l=h^{\prime}p^{s-1}$ we find
$\left.\dfrac{1}{(z^{p},q^{p})_{h^{\prime}p^{s-1}}}\right|_{q=\zeta}=\dfrac{1}{(1-z^{p^{m}})^{h^{\prime}p^{s-m}}}=\sum\limits_{i=0}^{\infty}\,\left.\dfrac{[q^{pi},q^{p}]_{h^{\prime}p^{s-1}}}{[1,q^{p}]_{h^{\prime}p^{s-1}}}\right|_{q=\zeta}z^{i}$
Comparing the coefficients we conclude
Lemma A.1.
Let $\zeta\neq 1$ be a unit $\zeta^{p^{s}}=1$ and let $h^{\prime}\in\mathbb{N}$ then for any $i\in\mathbb{N}$ we have
(A.6) $\left.\dfrac{[q^{i},q]_{h^{\prime}p^{s}}}{[1,q]_{h^{\prime}p^{s}}}\right|_{q=\zeta}=\left.\dfrac{[q^{pi},q^{p}]_{h^{\prime}p^{s-1}}}{[1,q^{p}]_{h^{\prime}p^{s-1}}}\right|_{q=\zeta}$
Coefficients (A.6) are equal to 00 unless $p\mid i$.
Lemma A.2.
a) If $h\in\mathbb{Z}$ then the hypergeometric function $F(h,\boldsymbol{a},q,z)$ (2.3) does not have poles at $q$ given by roots of unity.
b) Let $k\in\mathbb{Z}$ and $\zeta$ be a primitive unit root of order $m$ then:
$\left.F(mk,\boldsymbol{a},q,z)\right|_{q=\zeta}=\dfrac{1}{(1-z^{m})^{k}}$
Proof.
The proof follows directly from the definition of $q$-hypergeometric series. ∎
A.3. $p$-adic $q$-deformations
We recall the original definitions of the $p$-adic gamma and $q$-gamma functions:
Theorem A.3 (Koblitz, [Ko80]).
For $q=1+t$ with $|t|_{p}<1$ the function defined on $n\in\mathbb{N}$ by $\Gamma_{p,q}(n)=(-1)^{n}\prod\limits_{{0<i<n}\atop{p\nmid i}}\,[i]_{q}$ extends to a continuous function on $\mathbb{Z}_{p}$.
In the limit $t\to 0$ we have $\lim\limits_{q\to 1}\,\Gamma_{p,q}=\Gamma_{p}$ where $\Gamma_{p}$ denotes the $p$-adic gamma function of Morita [Mo75], which is the extension to $\mathbb{Z}_{p}$ of $\Gamma_{p}(n)=(-1)^{n}\prod\limits_{{0<i<n}\atop{p\nmid i}}\,n.$
Lemma A.4.
For $q=1+t$ with $|t|_{p}<1$, the function of $n,h\in\mathbb{N}$ given by
(A.7) $f(h,n)=\dfrac{\dfrac{(q^{ph},q)_{\infty}}{(q,q)_{\infty}}\dfrac{(q^{pn+1},q)_{\infty}}{(q^{pn+ph},q)_{\infty}}}{\dfrac{(q^{ph},q^{p})_{\infty}}{(q^{p},q^{p})_{\infty}}\dfrac{(q^{pn+p},q^{p})_{\infty}}{(q^{pn+ph},q^{p})_{\infty}}}$
extends to $\mathbb{Z}_{p}$ values of the arguments by $f(h,n)=\dfrac{\Gamma_{p,q}(pn+ph)}{\Gamma_{p,q}(pn)\,\Gamma_{p,q}(ph)}$.
Proof.
If $n$ and $h$ are two natural numbers then
$\dfrac{(q^{ph},q)_{\infty}}{(q,q)_{\infty}}\dfrac{(q^{pn+1},q)_{\infty}}{(q^{pn+ph},q)_{\infty}}=\binom{pn+ph-1}{pn}_{\!\!q}$
is a $q$-binomial coefficient (A.1). Similarly
$\dfrac{(q^{ph},q^{p})_{\infty}}{(q^{p},q^{p})_{\infty}}\dfrac{(q^{pn+p},q^{p})_{\infty}}{(q^{pn+ph},q^{p})_{\infty}}=\binom{n+h-1}{n}_{\!\!q^{p}}.$
For the ratio of these factors we find:
$\dfrac{\binom{pn+ph-1}{pn}_{\!\!q}}{\binom{n+h-1}{n}_{\!\!q^{p}}}=\dfrac{\prod\limits_{{0<i<pn+ph}\atop{p\nmid i}}\,[i]_{q}}{(\prod\limits_{{0<i<pn}\atop{p\nmid i}}\,[i]_{q})(\prod\limits_{{0<i<ph}\atop{p\nmid i}}\,[i]_{q})}=\dfrac{\Gamma_{p,q}(pn+ph)}{\Gamma_{p,q}(pn)\,\Gamma_{p,q}(ph)}$
By Theorem A.3 this function extends to $\mathbb{Z}_{p}$-values of $h$ and $n$. ∎
Let us recall that the $p$-adic unit roots $q\in\Omega$ of order $p^{s}$, are of the form $q=1+t$ with $|t|_{p}=p^{-\frac{1}{p^{s-1}(p-1)}}<1$, [Ko84]. By Theorem A.3 we can consider the values of $p$-adic $q$-gamma functions at such $q$. The following Lemma will be useful in the next section.
Lemma A.5.
let $x,y\in\mathbb{Z}_{p}$ and $q\neq 1$ be a $p^{s}$-th unit root, then
(A.8) $\Gamma_{p,q}(x+yp^{s})=\Gamma_{p,q}(x)\Gamma_{p,q}(yp^{s})$
Proof.
First we note that for $x\in\mathbb{Z}_{p}$ and $q\neq 1$ given by a $p^{s}$-th unit root we have:
(A.9) $\Gamma_{p,q}(x+p^{s})=\Gamma_{p,q}(x)\Gamma_{p,q}(p^{s})$
Indeed, for $x\in\mathbb{N}$ the identify (A.9) follows directly from the definition of $(p,q)$-Gamma function:
$\Gamma_{p,q}(x)=(-1)^{n}\prod\limits_{{0<i<x}\atop{p\nmid i}}\,[i]_{q}$
which extends to $x\in\mathbb{Z}_{p}$ as in Theorem A.3. Next, for $y\in\mathbb{N}$ (A.8) follows by applying (A.9) $y$-times. Then both sides of (A.8) extend to $y\in\mathbb{Z}_{p}$ as in Theorem A.3 again. ∎
A.4.
We can write
$\Gamma_{p,q}(p)=-\dfrac{1-q}{1-q}\dfrac{1-q^{2}}{1-q}\cdots\dfrac{1-q^{p-1}}{1-q}=-\lim\limits_{w\rightarrow 1}\,\dfrac{1-w}{1-w}\dfrac{1-wq}{1-q}\cdots\dfrac{1-wq^{p-1}}{1-q}$
where the limit is for regularization purposes. Now let us assume that $q$ is a primitive $p$-th root of unity, $q^{p}=1$ then from above formula we obtain:
$\Gamma_{p,q}(p)=\dfrac{-1}{(1-q)^{p-1}}\lim\limits_{w\rightarrow 1}\,\dfrac{1-w^{p}}{1-w}=\dfrac{-p}{(1-q)^{p-1}}$
Using same logic we also find for $m\in\mathbb{N}$:
$\Gamma_{p,q}(mp)=\Gamma_{p,q}(p)^{m}=\Big{(}\dfrac{-p}{(1-q)^{p-1}}\Big{)}^{m}$
For primitive $p^{s}$-th we can use a similar trick. First we write:
$\Gamma_{p,q}(p^{s})=-\prod\limits_{{0<i<p^{s}}\atop{p\nmid i}}\,\dfrac{1-q^{i}}{1-q}=\dfrac{-1}{(1-q)^{p^{s-1}(p-1)}}\lim\limits_{w\to 1}\,\dfrac{\prod\limits_{{0<i<p^{s}}}(1-wq^{i})}{\prod\limits_{{0<i<p^{s-1}}}(1-wq^{ip})}$
Assume $q$ is a primitive $p^{s}$-th root of unity, $q^{p^{s}}=1$, then the previous expression gives:
$\Gamma_{p,q}(p^{s})=\dfrac{-1}{(1-q)^{p^{s-1}(p-1)}}\lim\limits_{w\to 1}\dfrac{1-w^{p^{s}}}{1-w^{p^{s-1}}}=\dfrac{-p}{(1-q)^{p^{s-1}(p-1)}}$
In general we conclude
Proposition A.6.
Let $q$ be a primitive $p^{s}$-th unit root and $m\in\mathbb{N}$ then
$\Gamma_{p,q}(mp^{s})=\Big{(}\dfrac{-p}{(1-q)^{p^{s-1}(p-1)}}\Big{)}^{m}$
Note also that a primitive $p^{s}$-th unit root has the form $q=1+t$ with $|t|_{p}=p^{-\frac{1}{p^{s-1}(p-1)}}$. Thus we see that $|\Gamma_{p,q}(p^{s})|_{p}=1$.